### v2

# Learning Hierarchical Sparse Transform Coding of 3DGS

Abstract
3D Gaussian Splatting (3DGS) supports fast, high quality, novel view synthesis but has a heavy memory footprint, making the compression of its model crucial. Current state-of-the-art (SOTA) 3DGS compression methods adopt an anchor-based architecture that pairs the Scaffold-GS representation with conditional entropy coding. However, these methods forego the analysis-synthesis transform, a vital mechanism in visual data compression. As a result, redundancy remains intact in the signal and its removal is left to the entropy coder, which computationally overburdens the entropy coding module, increasing coding latency. Even with added complexity thorough redundancy removal is a task unsuited to an entropy coder. To fix this critical omission, we introduce a Sparsity-guided Hierarchical Transform Coding (SHTC) method, the first study on the end-to-end learned neural transform coding of 3DGS. SHTC applies KLT to decorrelate intra-anchor attributes, followed by quantization and entropy coding, and then compresses KLT residuals with a low-complexity, scene-adaptive neural transform. Aided by the sparsity prior and deep unfolding technique, the learned transform uses only a few trainable parameters, reducing the memory usage. Overall, SHTC achieves an appreciably improved R–D performance and at the same time higher decoding speed over SOTA. Its prior-guided, parameter-efficient design may also inspire low-complexity neural image and video codecs. Our code will be released at here.
Figure 1: Illustration of Scaffold-GS and anchor-based compression methods.
1 Introduction
In recent years, 3D Gaussian Splatting (3DGS) [kerbl20233d] has advanced computer vision and graphics by enabling real-time, high-quality, novel-view synthesis. The cost is substantial storage and bandwidth overheads incurred by storing a very large number of Gaussian primitives, making 3DGS compression an important research topic. Current state-of-the-art (SOTA) methods follow an anchor-based paradigm that pairs Scaffold-GS [lu2024scaffold] with context-based entropy coding to exploit spatial correlations and improve compression ratios [chen2024hac, wang2024contextgs, liu2024hemgs, zhan2025catdgs, chen2025hac++]. However, these methods forego end-to-end optimized analysis-synthesis transforms that remove redundancy before entropy coding. Without such an analysis transform, high-dimensional correlations persist and the inherent sparsity is left underutilized, forcing complex causal entropy models to compensate for the omission. Yet even large entropy models struggle to accurately capture inter-token dependencies, leading to suboptimal rate–distortion (R–D) performance. Additionally, because the 3DGS compression system operates in the “compress once, decompress many times” mode, decoding speed is critical; using complex entropy models inevitably increases decoding latency and degrades the client-side user experience.
To address aforementioned omission and drawbacks, we integrate an end-to-end optimized analysis–synthesis transform into the 3DGS compression pipeline, thereby removing the most of redundancy before entropy coding. Despite the promise of this paradigm and its maturity in neural image and video compression [balle2018variational, minnen2018joint, cheng2020learned, li2024neural], applying it to 3DGS compression requires overcoming two domain-specific challenges. First, point clouds in 3DGS are unstructured: unlike pixels on a grid, neighboring Gaussians (or anchors) cannot be indexed directly. While K-nearest neighbors (KNN) provides a way to define local neighborhoods on such irregular data, the spatial locations of Gaussians (or anchors) are variables in the R–D optimization and change during training, making per-iteration recomputation of KNN neighborhoods computationally prohibitive. Second, because each scene uses a dedicated transform whose parameters must be transmitted in the bitstream, the parameter budget must remain small; achieving strong performance under this constraint requires careful architectural design.
To tackle these challenges, we propose sparsity-guided hierarchical transform coding (SHTC), which (i) avoids spatial KNN by restricting transforms to channel-wise operations for computational efficiency and (ii) uses a two-layer architecture for parameter efficiency. Layer 1 applies the Karhunen-Loève Transform (KLT) for decorrelation and energy compaction, yielding channels with uneven energy distribution. Transmitting only the principal coefficients causes information loss and limits the achievable rendering quality, whereas coding all coefficients is rate-inefficient. To strike a balance between these extremes, we introduce a neural refinement module in Layer 2, which serves as a refinement layer. It compensates for truncation-induced information loss by transmitting the KLT residual with modest additional rate, using a more flexible strategy than direct coefficient coding. Because the residual components are predominantly small (with many values near zero), we model them as sparse. For such sparse signals, compressed sensing theory shows that the full signal can be reliably recovered from a small number of linear measurements rather than coding all coefficients directly [donoho2006compressed]. Motivated by this, we design the refinement layer following compressed-sensing principles. At the encoder, the residuals are projected into a low-dimensional latent space via a learnable linear transform that acts as a measurement matrix, and at the decoder they are reconstructed by unfolding the classical Iterative Shrinkage–Thresholding Algorithm (ISTA) [zibulevsky2010l1] into a sequence of network layers [gregor2010learning, zhang2020deep, zhang2018ista]. This sparse-recovery-inspired design enables accurate residual reconstruction with low parameter cost.
In conclusion, our key contribution is the introduction of the first end-to-end optimized transform coding framework for 3DGS compression, delivering substantial R-D gains with minimal additional parameters and runtime overhead. Across three standard datasets (Mip-NeRF 360 [barron2022mip], Tanks&Temples [knapitsch2017tanks], and DeepBlending [hedman2018deep]), our method reduces the memory footprint by $123.5\times$, $216.7\times$, and $705.3\times$ relative to vanilla 3DGS, respectively, while matching its rendering quality. In addition, our parameter-efficient design strategy, which combines a sparsity prior with deep unfolding, may offer a potential blueprint for developing low-complexity neural image and video codecs.
(a) HAC
(b) HAC++
(c) KLT coefficients
(d) HAC
(e) HAC++
(f) KLT coefficients
Figure 2: Comparison of inter-channel correlation (visualized as absolute Pearson correlation matrices) and energy compaction in the anchor features of HAC/HAC++ and their KLT coefficients, using the ‘playroom’ scene as an example.
2 Related Work
Following the taxonomy of [ali2025compression], we group existing 3DGS compression methods into (i) unstructured approaches, which employ model compression techniques such as pruning [ali2024elmgs, ali2024trimming, fan2024lightgaussian, niemeyer2024radsplat, lee2024compact, hanson2025pup] and quantization [girish2024eagles, navaneet2024compgs, wang2024end] for compression, and (ii) structured approaches [morgenstern2024compact, lee2025compression, lu2024scaffold], which exploit the spatial relationship among Gaussian primitives for compression. Because anchor-based designs are especially prevalent within structured approaches and our method builds on this architecture, we dedicate this section to anchor-based methods. Due to space limitations, a more detailed literature review is provided in the supplementary material.
Among anchor-based 3DGS compression methods, Scaffold-GS [lu2024scaffold] is the representative baseline. Scaffold-GS introduces a hierarchical structure during training: a sparse set of anchors serve as reference points from which a dense set of Gaussians are generated. As shown on the left side of Fig. 1, each anchor is parameterized by the position $\mathbf{p}$, latent feature $\mathbf{f}$, scaling factor $\mathbf{l}$, and $K$ learnable offsets $\{\mathbf{O}_{i}\}_{i=1}^{K}$. The scaling factor $\mathbf{l}$ together with the offsets determine the spatial distribution of the generated Gaussians. For each generated Gaussian, the view-dependent color, opacity, and covariance (parameterized by quaternion + scale) are predicted from the anchor’s latent feature through lightweight MLPs.
Although Scaffold-GS is more compact than vanilla 3DGS, it is essentially a 3DGS compaction scheme: anchors and their attributes are still stored as raw 32-bit floats without quantization or entropy coding, which inherently limits the achievable compression ratio. SOGS reduces the dimensionality of anchor features; however, this remains 3DGS compaction rather than true compression, so the gains are still limited [zhang2025sogs]. To address these limitations, recent work introduces context-based conditional entropy models that reduce statistical redundancy across anchor attributes, achieving significantly higher compression ratios. Among these methods, HAC was introduced early and serves as a foundational baseline. As illustrated in the right panel of Fig. 1, HAC uses multi-scale hash grids as a hyperprior [balle2018variational]. Given an anchor’s position, it queries these hash grids to obtain contextual features, which an MLP maps to the quantization step size $q_{s}$, the Gaussian likelihood parameters $(\boldsymbol{\mu},\boldsymbol{\sigma})$. HAC++ [chen2025hac++] enhances this design by (i) splitting each anchor feature vector into non-overlapping subgroups and encoding them sequentially with a channel-wise autoregressive model (ChARM) [minnen2020channel], and (ii) replacing a single Gaussian with a Gaussian mixture model to model more complex densities. CAT-3DGS [zhan2025catdgs] also adopts ChARM but substitutes hash-grid hyperpriors with triplanes. Both ContextGS [wang2024contextgs] and HEMGS [liu2024hemgs] follow a spatially sequential scheme: they use previously decoded anchors as conditioning context in a spatial autoregressive model (SARM) to encode the next anchors. These approaches represent the current state-of-the-art compression performance for 3DGS compression.
3 Research Motivation and Challenges
In this section, we analyze the importance and challenges of implementing end-to-end optimized transform coding in 3DGS compression.
3.1 Motivation
Decorrelation
Anchor-based compression methods, built on Scaffold-GS, achieve high compression ratios by quantizing and entropy-coding anchor attributes. Let $\mathbf{x}$ denote an anchor-attribute vector. With entropy coding, the ideal rate equals the source entropy $H_{q}=\mathbb{E}_{\mathbf{x}\sim q(\mathbf{x})}[-\log q(\mathbf{x})]$, but the ground truth distribution $q(\mathbf{x})$ is unknown. In practice, one can approximate the unknown distribution $q(\cdot)$ with a parametric model $p(\cdot)$ (often modeled as Gaussian), so the actual rate is the cross-entropy
$R=\mathbb{E}_{\mathbf{x}\sim q(\mathbf{x})}[-\log p(\mathbf{x})]=H_{q}+D_{KL}(q||p).$ (1)
where $D_{KL}(q||p)$ measures the extra bits caused by the mismatch between $q(\cdot)$ and $p(\cdot)$. Since anchor attributes are correlated, anchor-based methods use autoregressive models to exploit these correlations and reduce $D_{KL}$, thereby tightening the cross-entropy rate. For intra-vector correlation, $\mathbf{x}$ can be partitioned into groups of size $g$ and coded sequentially with the probability model $p(\mathbf{x}_{i:i+g}|\mathbf{x}_{<i})$ (e.g., ChARM as used in HAC++ [chen2025hac++] and CAT-3DGS [zhan2025catdgs]); for spatial correlation, anchors are coded causally with the probability model $p(\mathbf{x}^{(i)}|\{\mathbf{x}^{(j)}\}_{j\in\mathcal{C}_{i}})$ where $\mathcal{C}_{i}$ denotes the causal neighbors of anchor-$i$ (SARM as in ContextGS [wang2024contextgs] and HEMGS [liu2024hemgs]). While more complex entropy models can improve compression by exploiting data correlations during entropy coding, they also increase decoding latency. In 3DGS’s “compress once, decompress many times” pattern, decoding latency on the client side is critical; complex entropy models slow down decoding and degrade user experience. Moreover, even with more complex entropy models, fully capturing these dependencies remains difficult, leading to sub-optimal compression performance. These limitations arise from the absence of end-to-end optimized analysis–synthesis transforms. This motivates introducing such transforms into the compression pipeline: by decorrelating the signal before entropy coding, redundancy is reduced and much simpler, more efficient entropy models can be used. As shown in the first row of Fig. 2, HAC and HAC++ anchor features exhibit strong inter-channel correlations. Applying a channel-wise KLT largely removes these correlations, yielding decorrelated coefficients and thereby paving the way for the use of simpler entropy models.
Energy compaction
In the original signal domain, the relative importance of channels is unclear, hindering efficient rate allocation. After an analysis transform, energy concentrates in a small subset of coefficients, making their importance explicit and enabling more effective rate allocation. This effect is evident in the second row of Fig. 2, where KLT compacts most of the energy into the principal coefficients.
Applying an analysis transform before entropy coding decorrelates the data and compacts its energy, which benefits lossy compression, while a paired synthesis transform reconstructs the anchor attributes for rendering. When the transforms and entropy models are jointly optimized, this transform coding paradigm typically achieves superior R–D performance, as demonstrated in neural image compression [balle2016end, balle2018variational, minnen2018joint, cheng2020learned, he2022elic, zou2022devil, zhu2021transformer, liu2023learned, ali2024towards]. Motivated by these successes, extending this paradigm to 3DGS compression is a natural choice.
3.2 Challenges
Despite the promise of end-to-end optimized transform coding, to the best of our knowledge, no existing work has applied such transforms to 3DGS compression. Extending end-to-end optimized transform coding from neural image compression to 3DGS is, however, non-trivial due to two domain-specific challenges.
The first challenge stems from the irregular, unordered nature of 3DGS anchor points, which lack a natural neighborhood structure. Although neighborhoods can be defined via KNN, repeated neighbor searches in dense point clouds are computationally expensive. This issue is exacerbated by the fact that anchor positions are optimization variables that evolve during training, so the KNN graph must be rebuilt throughout optimization rather than constructed once, significantly increasing the overall training cost.
The second challenge arises from a fundamental difference between 3DGS compression and neural image/video codecs. In standard neural codecs, the analysis–synthesis transforms are trained once, shared across inputs, and omitted from the bitstream; consequently, large-capacity transforms can be used to pursue better R-D trade-offs. By contrast, 3DGS relies on scene-specific transforms whose parameters must be transmitted, imposing a strict parameter budget. Under this constraint, strong R–D performance hinges on architectural efficiency rather than sheer model size. Historically, early neural image codecs with relatively simple transforms [balle2016end, balle2018variational] failed to surpass traditional codecs, whereas later models with substantially more complex transforms did [cheng2020learned]. This evolution highlights the difficulty of designing neural transforms that are both parameter-efficient and capable of strong R-D performance.
These two challenges call for careful architectural design to balance R-D gains against additional computational and parameter costs. To address them, we propose SHTC, a transform-coding framework for 3DGS that is both computationally and parameter efficient, achieving strong R-D performance under a tight parameter budget. Technical details of SHTC are provided in Sec. 4.
4 SHTC: Sparsity-guided Hierarchical Transform Coding
[…]
4.1 Layer 1: Base Layer
Transform coding targets decorrelation and energy compaction; among orthonormal linear transforms, KLT is optimal for both. Under a tight parameter budget, we also find that an MLP-based transform lacks the capacity to surpass the KLT. Overall, the KLT offers an excellent cost–performance trade-off, achieving strong R-D performance with minimal compute and memory. We therefore adopt the KLT in Layer 1.
Let $\mathbf{f}\in\mathbb{R}^{N_{f}}$ denote the anchor feature, with global mean $\mathbf{m}\in\mathbb{R}^{N_{f}}$ and KLT basis $\mathbf{V}\in\mathbb{R}^{N_{f}\times N_{f}}$. The analysis transform is $\boldsymbol{\theta}=g_{a}(\mathbf{f})=\mathbf{V}^{\top}(\mathbf{f}-\mathbf{m})$. The KLT coefficients exhibit an uneven energy distribution. For the anchor feature, we retain only the top-$M$ principal coefficients $\boldsymbol{\theta}_{p}=\boldsymbol{\theta}_{1:M}$, which are quantized and entropy-coded. The base layer reconstruction is $\hat{\mathbf{f}}_{base}=\mathbf{V}_{:,1:M}\hat{\boldsymbol{\theta}}_{p}+\mathbf{m}$.
4.2 Layer 2: Refinement Layer via Sparsity-Aware Residual Coding
Discarding the remaining coefficients causes information loss. The artifacts introduced by this truncation become more pronounced at high rates, leading to a quality bottleneck. However, coding all coefficients is rate-inefficient. To mitigate this information loss while keeping the rate overhead modest, we introduce a neural refinement layer. This layer transmits the KLT residual to compensate for truncation. Because most of the signal energy is captured by the leading KLT coefficients, the residual is dominated by low-magnitude components and can be well approximated as (approximately) sparse. For sparse or compressible signals, compressed sensing theory shows that the full signal can be reliably reconstructed from a small number of linear measurements by solving a sparsity-regularized optimization problem [donoho2006compressed]. This insight motivates us to use a principled, sparsity-aware design for the refinement transforms. Instead of relying on a generic multilayer perceptron (MLP), we adopt a parameter-efficient transform inspired by compressed sensing and deep unfolding techniques [zhang2020deep, zhang2018ista], enabling flexible compression of the residual with low parameter cost.
4.2.1 Analysis Transform: Learned Linear Measurements of Residuals
Let $\mathbf{r}$ denote the KLT residual. We encode $\mathbf{r}$ using a learnable linear analysis transform $h_{a}(\cdot)$, which projects $\mathbf{r}$ into a lower-dimensional vector $\mathbf{y}\in\mathbb{R}^{N_{t}}$:
$\mathbf{y}=h_{a}(\mathbf{r})=\mathbf{A}\mathbf{r},$ (2)
where $\mathbf{A}$ is a learnable matrix. In the language of compressed sensing, $\mathbf{A}$ implements learned linear measurements of the residual. The measurement vector $\mathbf{y}$ is then quantized to $\hat{\mathbf{y}}$ and entropy coded.
4.2.2 Synthesis Transform: Sparsity-Regularized Residual Reconstruction
Given the decoded quantized measurements $\hat{\mathbf{y}}$, the synthesis transform $h_{s}(\cdot)$ reconstructs an estimate $\hat{\mathbf{r}}$ of the residual:
[…]
Rather than designing $h_{s}(\cdot)$ as a black-box neural network, we adopt an interpretable formulation that casts residual reconstruction as a sparsity-regularized inverse problem. Assuming that $\mathbf{r}$ admits a sparse representation under a learned dictionary $\mathbf{D}$, i.e., $\mathbf{r}\approx\mathbf{D}\boldsymbol{\beta}$ with sparse coefficients $\boldsymbol{\beta}$, the residual reconstruction can be formulated as the following optimization problem:
$\tilde{\boldsymbol{\beta}}=\arg\min_{\boldsymbol{\beta}}\frac{1}{2}\big\|\hat{\mathbf{y}}-\mathbf{A}\mathbf{D}\boldsymbol{\beta}\big\|_{2}^{2}+\gamma\big\|\boldsymbol{\beta}\big\|_{1}.$ (4)
where the estimation of residual is recovered as $\tilde{\mathbf{r}}=\mathbf{D}\tilde{\boldsymbol{\beta}}$.
This inverse problem can be addressed using a variety of well-established optimization algorithms. A classical choice is the Iterative Shrinkage-Thresholding Algorithm (ISTA) [zibulevsky2010l1]. At the $k$-th iteration, ISTA updates the current estimate by performing a gradient descent step on the quadratic data fidelity term, followed by a soft-thresholding operation to handle the non-smooth $\ell_{1}$ regularization term. The update rule is given by:
$)\boldsymbol{\beta}^{(k+1)}=\mathcal{S}_{\tau}\big(\boldsymbol{\beta}^{(k)}-\eta\mathbf{D}^{T}\mathbf{A}^{T}(\mathbf{A}\mathbf{D}\boldsymbol{\beta}^{(k)}-\hat{\mathbf{y}}$$\big)$ (5)
where $\eta$ denotes the step size, and $\mathcal{S}_{\tau}(\cdot)$ is the element-wise soft-thresholding function, defined as $\mathcal{S}_{\tau}(z)=\mathrm{sign}(z)\mathrm{max}(|z|-\tau,0)$. Directly applying ISTA, however, can be computationally expensive, as it typically requires many iterations and manual tuning of the hyperparameters $\tau$ and $\eta$. To address this, we adopt a deep unfolding approach [gregor2010learning, zhang2018ista, zhang2020deep] that maps a large number of ISTA iterations to a small number of learnable network layers, significantly reducing inference time while preserving interpretability and the sparsity prior. Each layer corresponds to one unfolded ISTA iteration, with layer-wise learnable step sizes and thresholds. In particular, $\tau$ and $\eta$ are learned separately for each channel in each layer. This design allows the synthesis transform to embed a learnable sparse structure, enabling efficient residual reconstruction from a compact representation with minimal overhead.
Concretely, the synthesis transform $h_{s}(\cdot)$ takes $\hat{\mathbf{y}}$ as input and initializes the sparse coefficient vector $\boldsymbol{\beta}^{(0)}$. Each unfolded ISTA layer then progressively updates the coefficients as
$\boldsymbol{\beta}^{(0)}\;\rightarrow\;\boldsymbol{\beta}^{(1)}\;\rightarrow\;\cdots\;\rightarrow\;\boldsymbol{\beta}^{(N_{s})},$ (6)
where $N_{s}$ is the number of ISTA layers. The final predicted residual is recovered as
$\hat{\mathbf{r}}=\mathbf{D}\boldsymbol{\beta}^{(N_{s})}.$ (7)
Finally, the predicted residual $\hat{\mathbf{r}}$ is combined with the base layer reconstruction to form the final anchor feature for rendering.
4.3 Implementation details
There are four types of objects that we need to compress:
  • Anchor positions: following HAC++ [chen2025hac++], we compress anchor coordinates using MPEG-GPCC [graziosi2020overview].
  • Features $\mathbf{f}\in\mathbb{R}^{50}$: we compress feature vectors with the full two-layer SHTC. The base layer encodes $\mathbf{f}$ into a low-dimensional representation $\boldsymbol{\theta}_{p}$ with $M=15$. The refinement layer then encodes the KLT residual $\mathbf{r}$ into a measurement vector $\mathbf{y}$ of dimension $N_{s}=15$.
  • Scalings $\mathbf{l}\in\mathbb{R}^{6}$: for the scaling vectors, we apply only the first SHTC layer (KLT) without dimensionality reduction. Because $\mathbf{l}$ is low-dimensional and the reconstruction quality is highly sensitive to it, truncating KLT coefficients in the transform domain severely degrades performance. Instead, for the KLT coefficients $\mathbf{l}_{t}\in\mathbb{R}^{6}$ we adopt a monotonically increasing, channel-wise quantization schedule:
    $q_{s}^{(i)}=q_{s}\exp(\alpha i),\quad i\in\{0,\ldots,5\},$ (8)
    where $\alpha$ is a learnable scalar and $q_{s}$ is the per-anchor base quantization step predicted by HAC. This quantization schedule acts as a remedial measure when dimensionality reduction cannot be applied: it assigns larger steps to lower-energy components, improving the rate–distortion trade-off while avoiding truncation loss.
  • Offsets $\{\mathbf{O}_{i}\}_{i=1}^{K}$: for offset vectors, we retain the original treatment in HAC. Due to offset masking, offset vectors have variable dimensionality across anchors, which makes a fixed transform impractical.
As described in Sec. 2 and illustrated in Fig. 1, HAC uses multi-scale hash grids to parameterize its entropy model. The original HAC queries multi-scale hash grids to predict the quantization step $q_{s}$ and Gaussian likelihood parameters $(\boldsymbol{\mu},\boldsymbol{\sigma})$ for three types of anchor attributes: $\mathbf{f}$, $\mathbf{l}$, and $\{\mathbf{O}_{i}\}_{i=1}^{K}$. After integrating SHTC into HAC, we slightly modify the entropy coding process. It still queries multi-scale hash grids, but now predicts $q_{s}$ and $(\boldsymbol{\mu},\boldsymbol{\sigma})$ for the transformed vectors $\boldsymbol{\theta}_{p}$, $\mathbf{y}$, and $\mathbf{l}_{t}$, as well as for the offset vectors $\{\mathbf{O}_{i}\}_{i=1}^{K}$. The remaining quantization and coding procedures are the same as in the original HAC.
Inspired by color-space transforms in traditional codecs, we replace the original pixel-wise $\ell_{1}$ loss between the rendered image $\hat{I}$ and ground truth $I$ with a YCbCr-based loss, $\mathcal{L}_{\mathrm{YCbCr}}$, which computes reconstruction errors in the YCbCr color space. Because the human visual system is more sensitive to luminance than to chrominance variations, we assign a higher weight to the Y component and add a sharpness term on Y together with a smoothness term on Cb/Cr to preserve structure while suppressing color noise. This loss regularizes the compressed 3DGS representation, yielding fewer luminance errors and lower perceptual distortion at a given bitrate and thereby modestly improving R–D performance. The full definition of $\mathcal{L}_{\mathrm{YCbCr}}$ is provided in the supplementary material.
The overall training loss is defined as
$\mathcal{L}=\mathcal{L}_{d}+\lambda\mathcal{L}_{\mathrm{b-rate}}+\lambda_{o}\mathcal{L}_{\mathrm{reg}}+\lambda_{e}\ell_{1}(\mathbf{r},\hat{\mathbf{r}})+\lambda_{r}\mathcal{L}_{\mathrm{r-rate}}$ (9)
[…]
5 Experimental results
5.1 Experimental setup
Dataset
We primarily evaluate the R-D performance on three commonly used large-scale real-scene datasets: DeepBlending [hedman2018deep], Mip-NeRF360 [barron2022mip], and Tanks&Temples [knapitsch2017tanks]. In particular, we assess all nine scenes from the Mip-NeRF360 dataset [barron2022mip]. Evaluations on additional datasets are provided in the supplementary material. These diverse datasets provide a comprehensive evaluation of the proposed SHTC framework.
Distortion metrics
We evaluate rendering distortion using three metrics: PSNR, SSIM [wang2004image], and LPIPS [zhang2018unreasonable].
Rate metrics
We adopt the BD-rate metric [bjontegaard2001calculation] to quantify memory savings by comparing the R–D curve of a compression method against that of a baseline. Specifically, BD-rate computes the average bitrate difference between the two R–D curves over a common distortion range. A negative BD-rate indicates that, on average, the method achieves a lower bitrate (measured in memory size) than the baseline at comparable distortion levels, with the absolute value (expressed as a percentage) quantifying the average bitrate reduction.
Training details
We follow the official training settings of HAC [chen2024hac] and HAC++ [chen2025hac++], and train each scene for 30k iterations. The Lagrange multiplier $\lambda$ is set to $\{0.002,0.004,0.008,0.015\}$ to obtain R–D curves that span a wide range of bitrates and distortions for a more comprehensive comparison. The gaps between successive $\lambda$ values are chosen to increase because the effect of $\lambda$ on the achieved bitrate is highly non-linear: especially in the low-bitrate regime, larger changes in $\lambda$ are required to produce clearly separated bitrates and thus more evenly distributed points on the R–D curve. We set $\lambda_{e}=0.03$ and $\lambda_{r}=\max(\lambda/4,0.001)$. All experiments are conducted on a server equipped with two RTX 4090 GPUs.
5.2 Comparison with Anchor-based Methods
5.2.1 Evaluating R-D performance
(a) Mip-NeRF360
(b) Tank&Temples
(c) DeepBlending
Figure 3: Comparison of our method with existing anchor-based 3DGS compression methods.
Fig. 3 shows the R-D curves of several anchor-based 3DGS compression methods. Across all three datasets, the curve corresponding to the proposed SHTC, labeled “Ours SHTC (Full)” in the legend, lies toward the upper-left region, indicating higher PSNR at a given size or, equivalently, smaller size at a given distortion. For PCGS [chen2025pcgs], we train the model with $\lambda\in\{0.004,0.008,0.015,0.025\}$ to span a wide distortion range, so that its R-D points exhibit sufficient PSNR variation and a modest amount of overlap with the other methods, making the differences in the bitrate required to reach a given PSNR clearly visible. For HEMGS [liu2024hemgs], which is not open-sourced, we plot the R-D points reported in their paper and conduct an approximate comparison. Using only half as many training iterations (30k vs 60k), our method performs on par with or better than HEMGS on Mip-NeRF360 and DeepBlending and appears competitive on Tanks&Temples.
Tab. 1 reports the corresponding BD-rate of SHTC relative to HAC++, CAT-3DGS, and ContextGS, quantifying the bitrate (memory footprint) savings at equal PSNR. On average over the three datasets, our method achieves BD-rate values of about -21.0%, -24.3%, and -43.6% relative to HAC++, CAT-3DGS, and ContextGS, respectively.
Taken together, these results indicate that combining end-to-end optimized transforms with a simple entropy model is more effective than solely increasing entropy-model complexity in anchor-based 3DGS compression. This highlights the importance of incorporating transform coding and demonstrates the strong performance of the proposed SHTC framework.
Dataset HAC++ CAT-3DGS (40k) ContextGS Mip-NeRF360 -20.81% -24.54% -49.36% Tank&Temples -22.55% -22.78% -39.17% DeepBlending -19.58% -25.53% -42.30%
Table 1: The BD-rate of SHTC relative to three baselines: HAC++ [chen2025hac++], CAT-3DGS [zhan2025catdgs] and ContextGS [wang2024contextgs]. Negative values indicate bitrate savings (better) compared with each baseline.
5.2.2 Evaluating Computational Cost
[…]
ContextGS HAC++ Ours ContextGS HAC++ Ours ContextGS HAC++ Ours ContextGS HAC++ Ours
Mip-NeRF360 106 128 146 32.73 8.48 5.90 33.32 13.86 8.86 -30.92% -56.60% -64.82%
Tank&Temples 130 158 177 29.37 5.63 3.91 29.05 8.81 5.80 -26.41% -44.87% -56.06%
DeepBlending 161 169 185 16.78 2.98 2.30 16.63 4.41 3.18 -41.09% -58.34% -64.56%
Table 2: Evaluating inference efficiency and compression effectiveness: complex entropy modeling alone vs. SHTC with a simple entropy model. From left to right, we report four metrics: rendering FPS, encoding time, decoding time, and BD-rate relative to the HAC baseline. All results are obtained from models trained with $\lambda=0.004$. The best result for each metric is highlighted in bold.
As shown in Tab. 2, our method not only achieves greater BD-rate gains relative to the HAC baseline than HAC++ and ContextGS, but also reduces both encoding and decoding latency. This efficiency gain stems from a paradigm shift: instead of relying on a highly complex entropy model to capture dependencies among high-dimensional anchor features, SHTC uses learned transforms to explicitly decorrelate anchor attributes before entropy coding. By shifting the decorrelation burden from the entropy model to the transform, SHTC produces coefficients for which a simple and fast entropy model is sufficient, yielding better overall compression efficiency and lower latency than relying on a highly complex entropy model alone.
In terms of training time, SHTC introduces a modest overhead. In our experiments, compared with HAC++, the training time per scene increases by at most about 8 minutes, while HAC++ itself already requires approximately 40–50 minutes to train. However, this additional cost is less critical in practice, because training is a one-time offline cost borne by 3DGS asset producers, whereas decoding is performed repeatedly on the user side: end users neither observe nor care about the training time, but they are directly sensitive to download and decoding latency. SHTC targets this more critical bottleneck by producing smaller bitstreams at comparable distortion and reducing decoding latency, making it particularly well suited for consumer-centric applications where fast access is essential. A detailed comparison of training times is provided in the supplementary material for reference.
In terms of parameter overhead, the transforms in SHTC introduce 5,093 additional parameters over the HAC baseline. However, when SHTC is integrated into HAC, the entropy model queries the hash grids and feeds an MLP that predicts probability parameters for low-dimensional transform coefficients instead of high-dimensional anchor features. This change reduces the size of this MLP. As a result, the integration of SHTC into HAC incurs a net overhead of only 1,154 parameters. In contrast, HAC++ increases the number of parameters by 45,400 relative to HAC. These results highlight the parameter efficiency of SHTC, which stems from a sparsity-aware transform design rather than simply adopting a larger black-box MLP.
5.2.3 Visual Comparison
Due to space limitations, visual comparisons are provided in the supplementary material.
(a) Mip-NeRF360
(b) Tank&Temples
(c) DeepBlending
Figure 4: Comparison of our method with several conceptually representative baslines.
5.3 Comparison with Other Methods
In addition to the comparisons with anchor-based compression methods using R–D curves and BD-rate, we also evaluate our method against a broad set of anchor-free and hybrid baselines [kerbl20233d, lu2024scaffold, lee2024compact, niedermayr2024compressed, girish2024eagles, fan2024lightgaussian, morgenstern2024compact, navaneet2024compgs, papantonakis2024reducing, wang2024end, liu2024compgs, chen2025fast, shin2025localityaware, tang2025neuralgs, hanson2025pup, xie2025sizegs, tian2025flexgaussian, lee2025optimized, tang2025feature, ma2025enhancing, morgenstern2024compact, lee2025compression, xie2024mesongs, yang2025hybridgs]. Because many of these methods only report one or two operating points on the R-D curve, they do not support a meaningful BD-rate computation; for these methods, we therefore provide a numerical comparison in tabular form in the supplementary material.
[…]
5.4 Ablation Study
As shown in Fig. 3, using only the base layer of SHTC, denoted as ‘Ours SHTC (BL)’, achieves slightly better or comparable R–D performance compared to HAC++ at low and medium bitrates. However, the improvements are modest overall, and SHTC (BL) fails to consistently surpass HAC++ in the high-rate region. This limited gain is mainly due to truncation-induced information loss in the base layer, which inevitably discards some useful information and constrains the achievable performance. In contrast, when we use the full SHTC to compress anchor attributes, which includes an additional neural refinement layer to compensate for truncation-induced information loss, the proposed method yields consistent gains over HAC++ across both low- and high-rate regions. At high bitrates, it also remains clearly superior in terms of R-D performance and does not suffer from the high-rate quality bottleneck observed with the base-layer-only variant.
The supplementary material provides a more detailed quantitative ablation study and a further analysis of the choice of transforms, and also discusses the limitations of our approach and directions for future work.
6 Conclusion
In this paper, we introduced SHTC, to our knowledge, the first end-to-end optimized transform coding framework for 3DGS compression, which achieves substantial R–D gains and improved inference efficiency with only minimal parameter overhead. By leveraging learned analysis–synthesis transforms for decorrelation instead of relying solely on highly complex entropy models, SHTC improves compression efficiency and reduces decoding latency, making it well suited to consumer-facing scenarios where fast transmission and decoding are critical. SHTC has the potential to inspire future work on 3DGS compression, for example by motivating more expressive transform architectures and spatially adaptive variants. More broadly, the idea of incorporating signal priors to improve the parameter efficiency of neural codecs may benefit the design of low-complexity neural image and video codecs.
[…]
Supplementary Material
(a) Synthetic-NeRF
(b) BungeeNeRF
Figure 5: Comparison of our method with anchor-based methods on two additional datasets, Synthetic-NeRF [mildenhall2021nerf] and BungeeNeRF [xiangli2022bungeenerf].
Figure 6: Visual comparison between HAC++ and our SHTC method on four scenes: ‘playroom’ (DeepBlending), ‘train’ (Tanks & Temples), ‘garden’ and ‘stump’ (Mip-NeRF360).
7 Overview of the Supplementary Material
This supplementary document provides additional technical details, extended experiments, and ablation studies that complement the main paper. Its content is organized as follows:
  • Sec. 8 compares our method with representative anchor-based methods on two additional datasets, Synthetic-NeRF [mildenhall2021nerf] and BungeeNeRF [xiangli2022bungeenerf].
  • Sec. 9 presents additional comparisons of computational cost between our approach and CAT-3DGS.
  • Sec. 10 shows visual comparisons.
  • Sec. 11 presents more detailed ablation studies.
  • Sec. 12 provides a more detailed discussion of our transform choices.
  • Sec. 13 presents additional comparisons in the training time.
  • Sec. 14 provides further comparisons with a set of widely used baselines and recent methods.
  • Sec. 15 summarizes implementation details.
    • Sec. 15.1 explains how we compute pixel-wise distortion in the YCbCr color space.
    • Sec. 15.2 details the selection of the $\lambda$ values used to generate the R-D curves in the main paper.
  • Sec. 16 discusses the limitations of our approach and outlines directions for future work.
  • Sec. 17 reviews additional related work to further contextualize our method.
8 Additional Results on the Synthetic-NeRF and BungeeNeRF Datasets
Fig. 5 compares the R-D curves of our method with several anchor-based 3DGS compression methods on the Synthetic-NeRF [mildenhall2021nerf] and BungeeNeRF [xiangli2022bungeenerf] datasets. As shown in the figure, our method consistently achieves better R-D performance, attaining a smaller memory footprint at comparable PSNR levels. Tab. 3 further reports the BD-rate gains of our method over three anchor-based baselines. Taken together, these additional experimental results further demonstrate the effectiveness of our approach and serve as a supplementary comparison to the experiments reported in the main text.
Dataset HAC++ CAT-3DGS (40k) HAC
Synthetic-NeRF -13.45% - -26.33%
BungeeNeRF -10.04% -39.90% -52.05%
Table 3: The BD-rate of our method relative to three baselines: HAC++ [chen2025hac++], CAT-3DGS [zhan2025catdgs] and HAC [chen2024hac]. Negative values indicate bitrate savings (better) compared with each baseline.
9 Additional Comparisons with CAT-3DGS
[…]
Table 4: Evaluating the effectiveness: CAT-3DGS vs. SHTC + simple entropy model. From left to right, we report four metrics: encoding time, decoding time, and BD-rate relative to the HAC baseline. The best result for each metric is highlighted in bold.
As a supplement to Table 2 in the main paper, Tab. 4 compares our method with CAT-3DGS. The evaluation metrics include encoding time, decoding time, and BD-rate relative to the HAC baseline. With HAC as the common baseline, our method achieves a larger reduction in memory footprint while requiring substantially less encoding and decoding time at inference. For CAT-3DGS [zhan2025catdgs], we encountered out-of-memory (OOM) errors during rendering when running the official implementation on our hardware, so the rendering FPS could not be measured. However, the training time, encoding time, and decoding time can still be measured before this OOM issue occurs. Consequently, we omit the rendering-FPS comparison for CAT-3DGS in the table.
10 Visual Comparison
We provide visual comparisons between our SHTC approach and HAC++ [chen2025hac++] in Fig. 6. For each example, we report the bitstream size (in MB) and the PSNR (in dB) of the shown crop. To illustrate how SHTC reduces rendering distortion under a smaller or comparable memory footprint, we evaluate four representative scenes: ‘playroom’ (DeepBlending), ‘train’ (Tanks & Temples), and ‘garden’ and ‘stump’ (Mip-NeRF360), which are shown from the first to the fourth row in Fig. 6. In the ‘playroom’ scene, our method better preserves the structure of the wall-mounted switch, which appears as an unrecognizable blur in the HAC++ reconstruction. In the ‘train’ scene, our method suppresses shading and dark banding artifacts along the painted stripes. In the ‘garden’ scene, our method removes the black floater near the table edge. In the ‘stump’ scene, our method effectively reduces floater artifacts around the tree stump.
(a) R-D curves for HAC-E, HAC++, and SHTC variants obtained by progressively adding the KLT-based base layer, scaling transform, YCbCr-based distortion, and the full two-layer SHTC.
(b) Comparison between our truncated-plus-refinement scheme (HAC-E + Full SHTC + YCbCr) and two variants that transmit all KLT coefficients without truncation.
Figure 7: Ablation Study of SHTC Variants on Mip-NeRF360.
11 Additional Ablation Study
11.1 Additional Quantitative Ablation Study
In the main paper, we only report the most basic configuration and the complete version of our method, mainly to let readers grasp the overall benefit of SHTC at a glance under the page limit. In this subsection, we present a more fine-grained quantitative ablation study to isolate and measure the contribution of each component that bridges the basic configuration and the full model.
To more accurately quantify the gains brought by the proposed transform modules, we construct an enhanced version of HAC, denoted by HAC-E. HAC-E is an intermediate variant between the vanilla HAC and HAC++. Following HAC++, HAC-E uses MPEG-GPCC to compress anchor positions and adopts the mask-aware rate estimator, but it still retains the simple entropy model of HAC and does not use complex context modeling. Starting from this common baseline, there are two ways to further improve compression performance:
  1. 1.
    replacing the simple entropy model with a highly complex one, which leads to HAC++;
  2. 2.
    keeping the simple, fast entropy model while introducing learned transforms to enhance R-D performance, which is our transform-coding paradigm.
Along the second route, we build several SHTC variants on top of HAC-E, and their BD-rate values over HAC-E directly quantify the gains of each component. Starting from HAC-E, we progressively add components:
  • HAC-E + BL(F15): we activate the most basic version of SHTC and use only its base layer (BL), implemented as a KLT, to compress anchor features. The dimensionality of the principal coefficients is set to 15.
  • HAC-E + BL(F15+S): we use the base layer to compress both anchor features and anchor scalings, as described in the main text.
  • HAC-E + BL(F15+S) + YCbCr: building on the previous variant, we replace the original pixel-wise $\ell_{1}$ distortion with the proposed YCbCr-space distortion $\mathcal{L}_{\mathrm{YCbCr}}$ in the training objective.
  • HAC-E + Full SHTC + YCbCr: finally, we use the full two-layer SHTC for anchor feature compression, so that the neural refinement layer compensates for truncation loss at low rate overhead, while the base layer is still used for anchor scaling compression. This corresponds to our full model.
We plot the R–D curves of these variants on the Mip-NeRF360 dataset in Fig. 7(a). Mip-NeRF360 contains diverse and challenging indoor and outdoor scenes, making it a suitable benchmark for ablation with reduced bias. From Fig. 7(a), we observe that introducing a KLT-based base layer for anchor feature compression (HAC-E + BL(F15)) already yields a clear gain over the HAC-E baseline and brings the performance close to HAC++. Extending the base layer to also compress anchor scalings (HAC-E + BL(F15+S)) and further switching to the YCbCr-based distortion metric (HAC-E + BL(F15+S) + YCbCr) provide additional, consistent improvements. The complete SHTC configuration trained with the YCbCr-based distortion (HAC-E + Full SHTC + YCbCr) achieves the best overall R-D trade-off among all variants.
Tab. 5 quantifies these effects by reporting BD-rate values over HAC-E on Mip-NeRF360. HAC++ is also included to facilitate a direct comparison between the two paradigms: using learned transforms with a simple entropy model versus relying solely on a highly complex entropy model. The results show that even the simplest transform-based variant (HAC-E + BL(F15)) already surpasses HAC++, and each additional component yields further BD-rate reductions, leading to a substantial $29.33\%$ reduction for HAC-E + Full SHTC + YCbCr.
BD-rate over HAC-E
HAC++ -10.93%
HAC-E+BL(F15) -13.30%
HAC-E+BL(F15+S) -16.59%
HAC-E+BL(F15+S)+YCbCr -20.07%
HAC-E+Full SHTC+YCbCr -29.33%
Table 5: BD-rate over HAC-E on the Mip-NeRF360 dataset for HAC++ and the proposed SHTC variants. Negative values indicate rate savings over HAC-E, and more negative values correspond to better R-D performance.
11.2 Comparison with No-Truncation Variants
Finally, we investigate whether directly transmitting all KLT coefficients can serve as a competitive alternative to our truncated-plus-refinement design. In Fig. 7(b), we first consider a variant that does not truncate the KLT coefficients of anchor features, denoted by HAC-E + BL(F50+S) + YCbCr. This configuration directly quantizes and entropy-codes all 50 KLT coefficients. Under the same rate-control parameter $\lambda$, it attains higher PSNR than the truncated-only baseline, but at a significantly higher bitrate, so it offers no improvement in overall R-D performance compared with simply truncating the coefficients.
Building on HAC-E + BL(F50+S) + YCbCr, we then introduce a monotonically increasing, channel-wise quantization schedule for the 50-dimensional KLT coefficients,
$q_{s}^{(i)}=q_{s}\exp(\alpha i),\qquad i\in\{0,\ldots,49\},$ (10)
where $\alpha$ is a learnable scalar and $q_{s}$ is the per-anchor base quantization step predicted by HAC. The idea is to reduce the bitrate cost of transmitting all coefficients by assigning larger quantization steps to less important channels. This variant, denoted by HAC-E + BL(F50+S) + YCbCr + CQ, yields a modest improvement over HAC-E + BL(F50+S) + YCbCr, but still falls short of our truncated-plus-refinement scheme in terms of R-D performance.
These results confirm that directly transmitting all KLT coefficients, even with a more sophisticated quantization schedule, is rate-inefficient and does not bring meaningful gains in R-D performance. In contrast, truncating the transform coefficients and compensating the lost information via a learned refinement layer provides a more effective and bitrate-efficient way to improve compression performance.
(a) HAC
(b) DCT coefficients
(c) KLT coefficients
(d) HAC
(e) DCT coefficients
(f) KLT coefficients
(g) Haar L-components
(h) Haar H-components
(i) Haar L-components
(j) Haar H-components
Figure 8: Inter-channel decorrelation and energy compaction across transforms. (a-c) Absolute inter-channel correlation matrices for the original data, DCT coefficients, and KLT coefficients. (d–f) Normalized energy per channel. (g-h) Absolute inter-channel correlation for the Haar transform’s two sub-bands. (i–j) Corresponding normalized energy distributions for Haar L and Haar H sub-bands.
12 Discussion of Transform Choices
12.1 Transform Choices for the Base Layer
In this section, we present a high-level comparison of transform design choices, supported by several analysis experiments.
Why Not DCT or DWT?
We considered several candidates for the base-layer transform, including the classical DCT and the Haar wavelet. To guide this choice, we compared their ability to decorrelate anchor channels and compact energy. Fig. 8 shows the absolute inter-channel correlation matrices and the normalized energy distributions of the original HAC features, the DCT coefficients, the Haar coefficients, and the KLT coefficients. DCT and Haar can reduce inter-channel correlation and slightly improve energy compactness, but their gains are limited: the correlation matrices still exhibit noticeable off-diagonal structure and the energy remains relatively spread out across channels. This behavior is expected because their bases are fixed and data-agnostic. In contrast, KLT yields an almost perfectly diagonal correlation matrix and concentrates most of the energy in the first few coefficients. This data-dependent transform is obtained as the eigenvectors of the sample covariance and is theoretically optimal for decorrelation and energy compaction among linear transforms. The main additional cost, relative to DCT and Haar, is storing the KLT basis. Even if the full basis must be transmitted, a 50-dimensional feature vector only requires a $50\times 50$ matrix (2,500 coefficients), which is negligible compared with the overall bitstream size.
Why Not an MLP-Based Transform?
Beyond classical transforms, we also compare our ‘HAC-E + KLT (F15)’ variant (the simplest version of SHTC) with a variant equipped with an MLP-based analysis and synthesis transform pair. Both networks are 5-layer MLPs, with a total of 21,930 parameters that must be transmitted as part of the bitstream. This comparison, evaluated on the Mip-NeRF360 dataset and shown in Fig. 9, reveals that the black-box MLP transform performs significantly worse in terms of the R-D trade-off than the simple KLT-based design. A likely reason is that, under a strict parameter budget, the MLP-based transform lacks sufficient expressive power to accurately reconstruct the input, leading to larger reconstruction errors and thus worse R-D performance.
For these reasons, we adopt KLT as a cost-effective choice for the base-layer transform. KLT provides strong decorrelation and energy compaction at negligible storage cost, while remaining simple and invertible. On top of this, we use a neural residual coder to compensate for the truncation loss of the KLT coefficients, which further improves the R-D performance over using KLT alone.
Figure 9: Comparison between ‘HAC-E + KLT (F15)’ and ‘HAC-E + MLP’ on the Mip-NeRF360 dataset.
12.2 Transform Choices for the Refinement Layer
When designing the neural refinement layer, a natural question arises: why not simply use a generic MLP? Our key observation is that the residual to be coded by the refinement layer (i.e., the KLT residual) can be reasonably modeled as a sparse signal. Classical compressed sensing theory shows that sparse signals can be recovered from a small number of linear measurements by solving a sparsity-regularized inverse problem. This perspective suggests a principled way to construct the refinement layer: instead of relying on an unconstrained black-box network, we explicitly exploit sparsity and build the module on top of a well-understood sparse reconstruction framework, so that each component has a clear interpretation.
In 3DGS compression, the parameters of transforms must be transmitted as part of the bitstream, so the parameter budget is limited. Under such constraints, a purely neural, black-box MLP may not have sufficient effective capacity to learn a good transform and its approximate inverse from data. In contrast, a sparsity-aware design embeds strong structure and prior knowledge directly into the architecture, so that even with relatively few parameters it can still achieve accurate reconstruction. This leads to much higher parameter efficiency than a generic MLP-based transform.
To verify this intuition, we conduct several offline experiments; here we present results on the bicycle scene of the Mip-NeRF360 dataset, where we compress the KLT residual in two different ways. The first variant uses a pure MLP-based transform with an MLP analysis and synthesis transform pair; both networks are 5-layer MLPs, resulting in a total of 21,930 parameters that must be transmitted in the bitstream. The second variant uses our sparsity-aware refinement layer, which contains only 5,093 parameters. We train both models for 3,000 iterations and measure the $\ell_{1}$ reconstruction loss of the KLT residual. As shown in Fig. 10, the sparsity-aware transform consistently drives the loss down to a much lower level than the MLP-based transform, despite using roughly four times fewer parameters. This early-stage verification supports our motivation: leveraging the sparsity prior and using a compressed-sensing-inspired architecture for residual coding yields a highly parameter-efficient refinement module. For this reason, we adopt the sparsity-aware design as our default choice for the neural refinement layer in SHTC.
Figure 10: Offline comparison of our sparsity-aware refinement layer and a black-box MLP-based transform on the bicycle scene of the Mip-NeRF360 dataset. The plot shows the $\ell_{1}$ reconstruction loss of the KLT residual versus training iteration.
13 Training time
Tab. 6 compares the training time of our method with three advanced anchor-based 3DGS compression methods that rely on complex entropy models. Although our method is not the fastest to train, its training time remains moderate: it requires more training time than HAC++, but consistently less than CAT-3DGS and ContextGS across all three benchmarks. Designing a compression system that simultaneously attains the best possible training efficiency, R-D performance, and inference speed is notoriously difficult; in practice, there is no free lunch. We argue that the resulting overhead in training time is acceptable because 3DGS compression follows a “compress once, decompress many times” paradigm. Training is a one-time offline cost borne by 3DGS asset producers, whereas decoding is performed repeatedly on the user side. End users are therefore insensitive to training time but directly affected by download and decoding latency. In this context, trading a modest increase in training time for improved R-D performance and faster decoding is both realistic and, in practice, a favorable design choice. Our method realizes this trade-off in practice, achieving improved R-D performance and lower decoding latency while keeping the training time within a reasonable range.
CAT-3DGS ContextGS HAC++ Ours
Mip-NeRF360 5725 3927 2735 3218
Tank&Temples 4034 2505 1880 2315
DeepBlending 3353 2320 1857 2270
Table 6: Training time (in seconds) of CAT-3DGS, ContextGS, HAC++, and our method on three datasets. The best and second-best results are highlighted in red and yellow cells, respectively.
Datasets Mip-NeRF360 Tank&Temples DeepBlending
[…]
3DGS [kerbl20233d] 27.46 0.812 0.222 750.9 23.69 0.844 0.178 431.0 29.42 0.899 0.247 663.9
Scaffold-GS [lu2024scaffold] 27.50 0.806 0.252 253.9 23.96 0.853 0.177 86.50 30.21 0.906 0.254 66.00
Compact3DGS [lee2024compact] 27.08 0.798 0.247 48.80 23.32 0.831 0.201 39.43 29.79 0.901 0.258 43.21
Compressed3D [niedermayr2024compressed] 26.98 0.801 0.238 28.80 23.32 0.832 0.194 17.28 29.38 0.898 0.253 25.30
EAGLES [girish2024eagles] 27.14 0.809 0.231 58.91 23.28 0.835 0.203 28.99 29.72 0.906 0.249 52.34
LightGaussian [fan2024lightgaussian] 27.00 0.799 0.249 44.54 22.83 0.822 0.242 22.43 27.01 0.872 0.308 33.94
Navaneet et al. [navaneet2024compgs] 27.12 0.806 0.240 19.33 23.44 0.838 0.198 12.50 29.90 0.907 0.251 13.50
Reduced3DGS [papantonakis2024reducing] 27.19 0.807 0.230 29.54 23.57 0.840 0.188 14.00 29.63 0.902 0.249 18.00
RDOGaussian [wang2024end] 27.05 0.802 0.239 23.46 23.34 0.835 0.195 12.03 29.63 0.902 0.252 18.00
PUP 3D-GS [hanson2025pup] 26.67 0.786 0.272 74.65 22.72 0.801 0.244 43.33 28.85 0.881 0.302 69.92
SizeGS [xie2025sizegs] 27.48 0.806 0.240 18.17 24.04 0.840 0.200 10.93 30.24 0.903 0.271 7.92
FlexGaussian [tian2025flexgaussian] 26.38 0.780 0.251 40.80 22.44 0.804 0.219 16.30 28.61 0.884 0.269 25.48
NeuralGS [tang2025neuralgs] 27.35 0.806 0.240 16.90 23.63 0.841 0.192 12.06 29.91 0.906 0.254 12.98
LocoGS [shin2025localityaware] 27.40 0.815 0.219 13.89 23.89 0.854 0.160 12.34 30.17 0.906 0.244 13.39
FCGS (low rate) [chen2025fast] 27.05 0.798 0.237 36.30 23.48 0.833 0.193 18.80 29.27 0.893 0.257 30.10
FCGS (high rate) [chen2025fast] 27.39 0.806 0.226 67.20 23.62 0.839 0.184 33.60 29.58 0.899 0.248 54.50
SOG [morgenstern2024compact] 26.56 0.791 0.241 16.70 23.15 0.828 0.198 9.30 29.12 0.892 0.270 5.70
CodecGS [lee2025compression] 27.30 0.810 0.236 9.78 23.63 0.842 0.192 7.46 29.82 0.907 0.251 8.62
MesonGS [xie2024mesongs] 26.99 0.796 0.247 27.16 23.32 0.837 0.193 16.99 29.51 0.901 0.251 24.76
CompGS [liu2024compgs] 27.26 0.803 0.239 16.50 23.70 0.837 0.208 9.60 29.69 0.901 0.279 8.77
FP-Net [ma2025enhancing] 27.46 0.801 0.249 11.08 24.13 0.847 0.192 6.34 30.14 0.906 0.274 3.46
FP-GS [tang2025feature] 27.53 0.806 0.240 12.82 24.32 0.811 0.185 7.32 30.05 0.906 0.266 3.91
HAC [chen2024hac] 27.53 0.807 0.238 15.26 24.04 0.846 0.187 8.10 29.98 0.902 0.269 4.35
Context-GS [wang2024contextgs] 27.62 0.808 0.237 12.68 24.20 0.852 0.184 7.05 30.11 0.907 0.258 3.45
HAC++ [chen2025hac++] 27.60 0.803 0.253 8.34 24.22 0.849 0.190 5.18 30.16 0.907 0.266 2.91
Ours 27.69 0.804 0.248 8.10 24.34 0.845 0.191 5.19 30.19 0.903 0.272 2.49
Table 7: Comparison of our SHTC framework with other 3DGS data compression methods, including 3DGS and Scaffold-GS, for reference. The best and second-best results are highlighted in red and yellow cells, respectively. The size values are measured in megabytes (MB).
14 Additional Comparisons with Other Methods
In addition to the comparisons with anchor-based compression methods using R–D curves and BD-rate, we also evaluate our method against a broad set of anchor-free and hybrid baselines. These include several commonly used 3DGS and compression baselines [kerbl20233d, lu2024scaffold, lee2024compact, niedermayr2024compressed, girish2024eagles, fan2024lightgaussian, morgenstern2024compact, navaneet2024compgs, papantonakis2024reducing, wang2024end, liu2024compgs], the feed-forward compression method FCGS [chen2025fast], and two neural-field-based 3DGS compression methods, LocoGS [shin2025localityaware] and NeuralGS [tang2025neuralgs]. We further consider several recent 3DGS compression methods [hanson2025pup, xie2025sizegs, tian2025flexgaussian, lee2025optimized], two HAC-based approaches that introduce prediction modules to improve HAC [tang2025feature, ma2025enhancing], and a group of methods that apply transform coding in a post-training stage for 3DGS compression [morgenstern2024compact, lee2025compression, xie2024mesongs, yang2025hybridgs].
Regarding SOGS [zhang2025sogs], we do not include it in our quantitative comparison for two reasons. First, to the best of our knowledge, there is no official implementation available, and the original paper does not report file sizes or distortion metrics on the datasets used in our experiments, which makes a fair and reproducible comparison difficult. Second, SOGS is essentially a 3DGS compaction method rather than a full compression scheme with an explicit bitstream and entropy coding. Based on the model statistics reported in [zhang2025sogs], we can only roughly infer an effective compression ratio of about $2\times$ over Scaffold-GS; this is merely a rough estimate, since the authors do not provide explicit compression ratios. In contrast, by employing the proposed SHTC framework, our method achieves substantially higher compression ratios over Scaffold-GS. For these reasons, we omit SOGS from the quantitative comparison.
Because many of the above methods report only one or two operating points on the R–D curve, a meaningful BD-rate computation is not feasible. For these methods, we therefore provide a numerical comparison in tabular form in Tab. 7. As shown in Tab. 7, our method achieves a better memory-quality trade-off than above approaches.
HAC ContextGS HAC++ Ours Mip-NeRF360 $\{0.002,0.004,0.008,0.015\}$ $\{0.002,0.004,0.008,0.015,0.025\}$ $\{0.002,0.004,0.008,0.015,0.025\}$ $\{0.002,0.004,0.008,0.015\}$ Tank&Temples $\{0.002,0.004,0.008,0.015,0.025\}$ $\{0.002,0.004,0.008,0.015,0.025,0.04\}$ $\{0.002,0.004,0.008,0.015,0.025\}$ $\{0.002,0.004,0.008,0.015\}$ DeepBlending $\{0.002,0.004,0.008,0.015,0.025\}$ $\{0.002,0.004,0.008,0.015,0.025,0.04\}$ $\{0.002,0.004,0.008,0.015,0.025,0.04\}$ $\{0.002,0.004,0.008,0.015,0.02,0.025\}$ Synthetic-NeRF $\{0.0005,0.001,0.002,0.003,0.004\}$ - $\{0.0005,0.001,0.002,0.003,0.004\}$ $\{0.0005,0.001,0.002,0.003,0.004\}$ BungeeNeRF $\{0.001,0.002,0.003,0.004,0.006\}$ $\{0.001,0.004\}$ $\{0.0005,0.001,0.002,0.003,0.004,0.006,0.008\}$ $\{0.001,0.002,0.003,0.004,0.006\}$
Table 8: Summary of $\lambda$ values used to generate R-D curves on different datasets.
15 Implementation Details
15.1 Pixel-wise Error in YCbCr Color Space
We propose a new distortion term for supervising the rendered images, inspired by the YCbCr color transform used in traditional image and video codecs. Because humans are more sensitive to luminance variations than to chrominance, we compute pixel-wise $\ell_{1}$ loss in the YCbCr space and place a higher weight on the Y component. To preserve fine detail, we add a high-frequency fidelity term on the Y channel by encouraging small differences between the Laplacian-filter response maps of $I_{Y}$ and $\hat{I}_{Y}$. To suppress color noise and artifacts, we further impose a total variation (TV) regularizer on the Cb and Cr channels. Together, the distortion term is defined as
$\mathcal{L}_{\mathrm{YCbCr}}=$ $\ell_{1}(I_{Y},\hat{I}_{Y})+\lambda_{c}\ell_{1}(I_{Cb},\hat{I}_{Cb})+\lambda_{c}\ell_{1}(I_{Cr},\hat{I}_{Cr})$
$+$ $\lambda_{H}\ell_{1}(\nabla^{2}I_{Y},\nabla^{2}\hat{I}_{Y})$ (11)
$+$ $\lambda_{TV}\mathrm{TV}(\hat{I}_{Cb})+\lambda_{TV}\mathrm{TV}(\hat{I}_{Cr}).$
This distortion term acts as a regularizer on the compressed 3DGS representation, encouraging it to render images with reduced luminance errors at a given bitrate, since luminance errors are perceptually more important than chrominance errors. This, in turn, reduces the overall distortion and empirically improves the R-D performance. In our implementation, we set $\lambda_{c}$, $\lambda_{H}$, and $\lambda_{TV}$ to 0.6, 0.15, and 0.1, respectively.
15.2 Details of $\lambda$ Selection for R-D Curves
In this subsection, we summarize in Tab. 8 the $\lambda$ values used to generate the R–D curves in our experiments. In addition to these values, we also clarify several special cases regarding how $\lambda$ and the corresponding R-D points are obtained for some baselines:
  • Main text. For the R-D curves in the main paper, we train the model with $\lambda\in\{0.002,0.004,0.008,0.015\}$, which yields a PSNR range of roughly 1 dB. The spacing between consecutive $\lambda$ value is gradually increased so as to cover a wide bitrate range while providing approximately uniform sampling of the R–D curves within that range. Because different models and datasets exhibit different sensitivity to $\lambda$, we additionally use larger $\lambda$ values for some baselines so that the PSNR range of each curve is comparable to that of the others.
  • CAT-3DGS. For CAT-3DGS [zhan2025catdgs], we encountered out-of-memory (OOM) errors during rendering when using the official implementation on our hardware. Consequently, we do not re-train or re-evaluate this method; instead, we directly use the R-D points reported in the original paper as its results in our plots.
  • Synthetic-NeRF. For experiments on the Synthetic-NeRF dataset, in order to obtain a quick verification of our method under comparable settings, we follow the $\lambda$ configurations recommended in the original HAC and HAC++. For the HAC, HAC++, and ContextGS baselines, we directly adopt the R-D numbers reported by their authors. Our own method is evaluated by running our implementation with the same $\lambda$ values.
  • BungeeNeRF. On the BungeeNeRF dataset, we again start from the $\lambda$ settings used in the original HAC, HAC++, and ContextGS papers for all methods, including ours. For these baselines, we use the R-D results reported by the authors at their original operating points. In addition, to obtain a PSNR range that is comparable to that of CAT-3DGS for a more reliable BD-rate computation, we further evaluate our method as well as HAC and HAC++ with one or two larger $\lambda$ values, thereby extending the PSNR variation range of these methods on BungeeNeRF.
  • PCGS. For PCGS, on both Synthetic-NeRF and BungeeNeRF, we found it difficult to obtain a sufficiently wide PSNR range for a stable BD-rate estimation by tuning $\lambda$. Therefore, we only plot the R-D points reported in the original paper and treat PCGS as a qualitative reference rather than including it in the quantitative BD-rate comparison.
16 Limitation and Future Work
In this work, SHTC adopts transforms that is shared across all spatial regions. This global sharing keeps the parameter count and bitstream overhead small and already yields strong R–D gains over existing methods, but it also limits the ability of the transform to adapt to local statistics. A natural extension to further improve performance is to introduce limited spatial adaptivity. For example, one could partition the anchors into blocks and learn a compact bank of transforms, transmitting a few index bits to select the best transform for a given anchor or block, analogous to intra prediction mode selection in video codecs. Since SHTC is highly parameter-efficient, the additional memory footprint of such a transform bank and the overhead of signaling the transform index would remain modest. As our primary goal in this work is to demonstrate the necessity and effectiveness of incorporating transform coding into the 3DGS training process, we leave the design and evaluation of these spatially adaptive variants for future work.
17 Supplementary Related Work
17.1 Unstructured 3DGS compression methods
3DGS has recently emgered as a pioneering approach for 3D reconstruction and representation, offering both high-quality and real-time rendering. Specifically, it represents a 3D scene as a collection of Gaussian primitives. Each Gaussian is parameterized by a position vector and a covariance matrix, which together define its position, shape and orientation in 3D space. Additionally, each Gaussian is associated with an opacity parameter and a set of Spherical Harmonics (SH) coefficients to model view-dependent colors. Through differentiable rendering techniques, all attributes of the Gaussians can be optimized, and the number of Gaussians can be progressively increased to minimize rendering distortion. However, the unconstrained Gaussian clone/splitting operation can generate millions of Gaussians, leading to the substantial burden on storage and bandwidth. This motivates researchers to explore 3DGS compression. Three observations motivate compressing 3D Gaussian Splatting (3DGS): (i) many Gaussians have negligible impact on rendering quality; (ii) many attributes tolerate reduced numerical precision; and (iii) only a subset of regions require high-order SH coefficients for sharp view-dependent effects. Taken together, these enable effective compression via pruning, SH distillation, and quantization.
Pruning
The densification process of 3DGS results in an explosion in the number of Gaussians. While this allows for the reconstruction of finer scene details, a substantial portion of the Gaussians is redundant, and removing them has little effect on visual quality. The challenge lies in identifying these unimportant Gaussians. [navaneet2024compgs] simply uses opacity as a criterion to remove transparent or nearly invisible Gaussians. [ali2024elmgs, ali2024trimming] further use both gradient magnitude and opacity levels as reference to identify removable Gaussians. Several subsequent studies assess each Gaussian’s contribution to ray color prediction and utilize this as the importance score [fan2024lightgaussian, girish2024eagles, niemeyer2024radsplat]. Lee et al. learn a binary mask to prune those unimportant Gaussians [lee2024compact], which is adopted by several subsequent works [wang2024end]. [hanson2025pup] propose a principled sensitivity pruning score by computing a second-order approximation of the reconstruction error with respect to the parameters of each Gaussian.
SH distillation
The largest portion of the memory footprint is used to store SH coefficients for modeling view-dependent color. However, a large fraction of the region consists of diffuse materials, which can be effectively modeled with view-independent color. This leads to significant waste in using high-degree SHs. To address this, [papantonakis2024reducing] assign each Gaussian an appropriate SH degree to reduce waste, while [fan2024lightgaussian] distill knowledge from a high-degree SH teacher model to a student model with truncated, lower-degree SHs.
Scalar quantization
Gaussian attributes are typically stored at 32-bit precision. However, opacity, scale, rotation, and SH coefficients are tolerant to small inaccuracies, suggesting that lower bit-depths can represent them with minimal performance drop. In light of this, [papantonakis2024reducing] propose a scalar quantization scheme that assigns each full-precision value to its nearest codeword in a K-means-derived codebook, storing only the corresponding index. [girish2024eagles] learn a latent representation, quantize it using a uniform scalar quantizer, and employ a decoder to predict Gaussian attributes from the quantized representation.
Vector quantization
Many Gaussians share similar parameters, allowing similar vectors to be mapped to a common code vector in a codebook. Instead of storing high-precision vectors, only compact indices are needed. Based on this insight, several works apply vector quantization to eliminate the redundancy within Gaussian attributes [navaneet2024compgs, fan2024lightgaussian, niedermayr2024compressed]. Furthermore, [wang2024end] employ entropy-constrained vector quantization, incorporating entropy coding into the quantization process. It assigns shorter codes to those code vectors that are used more frequently, effectively compressing the index stream.
17.2 Transform coding in NeRF/3DGS compression
Before the emergence of 3DGS, grid-based NeRF representations [muller2022instant, fridovich2023k, sun2022direct, chen2022tensorf] were considered a promising alternative to the classic NeRF [mildenhall2021nerf]. Several works have explored the use of transform coding to compress the feature grids in these representations [rho2023masked, li2024nerfcodec, pham2024neural]. Specifically, Rho et al. use a Discrete Wavelet Transform (DWT) to exploit transform-domain sparsity for compressing feature planes [rho2023masked]. Li et al. repurpose neural image codecs as learned nonlinear transforms: for each scene, they pre-train a plane-based NeRF, then scene-specifically fine-tune a subset of neural codec layers and encode the pretrained feature planes with the fine-tuned codec [li2024nerfcodec]. Pham et al. follow the same two-stage recipe but replace the off-the-shelf codec with lightweight nonlinear transforms designed from scratch [pham2024neural].
Compared with grid-based NeRFs, applying transform coding to 3DGS is more challenging because 3DGS data are irregular and unordered. To sidestep these challenges, prior work defers the application of transform coding until the post-training stage, once the 3DGS geometry is fixed. Two main strategies have been explored. First, some methods project 3DGS onto a set of feature planes [morgenstern2024compact] or adopt a triplane representation [chan2022efficient] to implicitly encode Gaussian attributes [lee2025compression], then compress the learned planes with standard image or video codecs. Second, other methods treat a pretrained 3DGS model as a point cloud and apply point-cloud attribute transforms such as Graph Fourier transform (GFT) [sandryhaila2013discrete] and Region-adaptive Hierarchical Transform (RAHT) [de2016compression] to compress the model [xie2024mesongs, yang2024benchmark, huang2025hierarchical, wang2025adaptive].
The above methods borrow transforms from images, video, or point clouds but apply fixed, linear transforms only after training. This decoupling blocks mutual adaptation between the transform and the 3DGS representation, often limiting R–D performance. In contrast, we propose an end-to-end transform-coding framework that jointly adapts the transform and the 3DGS representation, improving compression performance over post-training transform approaches.

### v3

@@
-Learning Hierarchical Sparse Transform Coding of 3DGS
+Learning Hierarchical Sparse Transform Coding for 3DGS Compression
 Abstract
-3D Gaussian Splatting (3DGS) supports fast, high quality, novel view synthesis but has a heavy memory footprint, making the compression of its model crucial.
-Current state-of-the-art (SOTA) 3DGS compression methods adopt an anchor-based architecture that pairs the Scaffold-GS representation with conditional entropy coding.
-However, these methods forego the analysis-synthesis transform, a vital mechanism in visual data compression.
-As a result, redundancy remains intact in the signal and its removal is left to the entropy coder, which computationally overburdens the entropy coding module, increasing coding latency.
-Even with added complexity thorough redundancy removal is a task unsuited to an entropy coder.
-To fix this critical omission, we introduce a Sparsity-guided Hierarchical Transform Coding (SHTC) method, the first study on the end-to-end learned neural transform coding of 3DGS.
-SHTC applies KLT to decorrelate intra-anchor attributes, followed by quantization and entropy coding, and then compresses KLT residuals with a low-complexity, scene-adaptive neural transform.
-Aided by the sparsity prior and deep unfolding technique, the learned transform uses only a few trainable parameters, reducing the memory usage.
-Overall, SHTC achieves an appreciably improved R–D performance and at the same time higher decoding speed over SOTA.
-Its prior-guided, parameter-efficient design may also inspire low-complexity neural image and video codecs.
-Our code will be released at here.
-Figure 1: Illustration of Scaffold-GS and anchor-based compression methods.
+Current 3DGS compression methods largely forego the neural analysis-synthesis transform, which is a crucial component in learned signal compression systems.
+As a result, redundancy removal is left solely to the entropy coder, overburdening the entropy coding module and reducing rate-distortion (R-D) performance.
+To fix this critical omission, we propose a training-time transform coding (TTC) method that adds the analysis-synthesis transform and optimizes it jointly with the 3DGS representation and entropy model.
+Concretely, we adopt a hierarchical design: a channel-wise KLT for decorrelation and energy compaction, followed by a sparsity-aware neural transform that reconstructs the KLT residuals with minimal parameter and computational overhead.
+Experiments show that our method delivers strong R–D performance with fast decoding, offering a favorable BD-rate–decoding-time trade-off over SOTA 3DGS compressors.
+Machine Learning, ICML
+Figure 1: Overview of 3DGS compression paradigms.
+Methods are categorized into four paradigms: (1) pruning and quantization, (2) anchor-based structure with adaptive entropy coding, (3) post-training transform coding (PTC) on a frozen 3DGS, and (4) training-time transform coding (TTC), introduced by us, which is the first to jointly optimize 3DGS, context model and the transform for improved R-D performance.
+Within TTC, we further propose SHTC, a lightweight transform architecture that delivers strong R-D performance with minimal parameter and computation overhead.
 1 Introduction
-In recent years, 3D Gaussian Splatting (3DGS) [kerbl20233d] has advanced computer vision and graphics by enabling real-time, high-quality, novel-view synthesis.
-The cost is substantial storage and bandwidth overheads incurred by storing a very large number of Gaussian primitives, making 3DGS compression an important research topic.
-Current state-of-the-art (SOTA) methods follow an anchor-based paradigm that pairs Scaffold-GS [lu2024scaffold] with context-based entropy coding to exploit spatial correlations and improve compression ratios [chen2024hac, wang2024contextgs, liu2024hemgs, zhan2025catdgs, chen2025hac++].
-However, these methods forego end-to-end optimized analysis-synthesis transforms that remove redundancy before entropy coding.
-Without such an analysis transform, high-dimensional correlations persist and the inherent sparsity is left underutilized, forcing complex causal entropy models to compensate for the omission.
-Yet even large entropy models struggle to accurately capture inter-token dependencies, leading to suboptimal rate–distortion (R–D) performance.
-Additionally, because the 3DGS compression system operates in the “compress once, decompress many times” mode, decoding speed is critical; using complex entropy models inevitably increases decoding latency and degrades the client-side user experience.
-To address aforementioned omission and drawbacks, we integrate an end-to-end optimized analysis–synthesis transform into the 3DGS compression pipeline, thereby removing the most of redundancy before entropy coding.
-Despite the promise of this paradigm and its maturity in neural image and video compression [balle2018variational, minnen2018joint, cheng2020learned, li2024neural], applying it to 3DGS compression requires overcoming two domain-specific challenges.
-First, point clouds in 3DGS are unstructured: unlike pixels on a grid, neighboring Gaussians (or anchors) cannot be indexed directly.
-While K-nearest neighbors (KNN) provides a way to define local neighborhoods on such irregular data, the spatial locations of Gaussians (or anchors) are variables in the R–D optimization and change during training, making per-iteration recomputation of KNN neighborhoods computationally prohibitive.
-Second, because each scene uses a dedicated transform whose parameters must be transmitted in the bitstream, the parameter budget must remain small; achieving strong performance under this constraint requires careful architectural design.
-To tackle these challenges, we propose sparsity-guided hierarchical transform coding (SHTC), which (i) avoids spatial KNN by restricting transforms to channel-wise operations for computational efficiency and (ii) uses a two-layer architecture for parameter efficiency.
-Layer 1 applies the Karhunen-Loève Transform (KLT) for decorrelation and energy compaction, yielding channels with uneven energy distribution.
-Transmitting only the principal coefficients causes information loss and limits the achievable rendering quality, whereas coding all coefficients is rate-inefficient.
-To strike a balance between these extremes, we introduce a neural refinement module in Layer 2, which serves as a refinement layer.
-It compensates for truncation-induced information loss by transmitting the KLT residual with modest additional rate, using a more flexible strategy than direct coefficient coding.
-Because the residual components are predominantly small (with many values near zero), we model them as sparse.
-For such sparse signals, compressed sensing theory shows that the full signal can be reliably recovered from a small number of linear measurements rather than coding all coefficients directly [donoho2006compressed].
-Motivated by this, we design the refinement layer following compressed-sensing principles.
-At the encoder, the residuals are projected into a low-dimensional latent space via a learnable linear transform that acts as a measurement matrix, and at the decoder they are reconstructed by unfolding the classical Iterative Shrinkage–Thresholding Algorithm (ISTA) [zibulevsky2010l1] into a sequence of network layers [gregor2010learning, zhang2020deep, zhang2018ista].
-This sparse-recovery-inspired design enables accurate residual reconstruction with low parameter cost.
-In conclusion, our key contribution is the introduction of the first end-to-end optimized transform coding framework for 3DGS compression, delivering substantial R-D gains with minimal additional parameters and runtime overhead.
-Across three standard datasets (Mip-NeRF 360 [barron2022mip], Tanks&Temples [knapitsch2017tanks], and DeepBlending [hedman2018deep]), our method reduces the memory footprint by $123.5\times$, $216.7\times$, and $705.3\times$ relative to vanilla 3DGS, respectively, while matching its rendering quality.
+3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) enables real-time, high-quality novel-view synthesis, but its many Gaussian primitives create significant storage and bandwidth overheads, motivating 3DGS compression.
+As summarized in Fig.
+1, existing 3DGS compression methods can be broadly grouped into unstructured approaches (e.g., pruning and quantization) and structured, entropy-coded pipelines.
+Within the structured category, most competitive compressors either adopt the anchor-based paradigm by pairing Scaffold-GS (Lu et al., 2024) with context-based entropy coding, or perform post-training transform coding (PTC) on a fixed 3DGS.
+In anchor-based methods, end-to-end analysis–synthesis transform coding is absent, so redundancy is not sufficiently removed before entropy coding.
+This leaves high-dimensional dependencies and sparsity underexploited, shifting the burden to increasingly complex entropy models (e.g., HAC++ (Chen et al., 2025c), ContextGS (Wang et al., 2024b)), which can still yield suboptimal rate-distortion (R–D) performance and higher decoding latency.
+Although a few prior works recognize the importance of transforms (Morgenstern et al., 2024; Lee et al., 2025b), they apply them only post-hoc to a fixed 3DGS, decoupling transform from 3DGS representation learning and preventing mutual adaptation under a unified R–D objective, which often limits compression gains.
+To address these limitations, we advocate a new paradigm, training-time transform coding (TTC), in which analysis–synthesis transforms are learned during training, allowing the 3DGS representation, entropy model, and transforms to be jointly optimized under a unified R-D objective.
+Unlike neural image/video compression where the transform is shared across images and incurs no per-image overhead, TTC uses scene-specific transforms that must be transmitted as part of the bitstream.
+From the minimum description length (MDL) perspective (Rissanen, 1978), TTC should ideally balance the transform cost and the coding cost of the transformed 3DGS attributes, i.e., $L(M)+L(D|M)$, where $L(M)$ denotes the number of bits to describe the scene-specific transform and $L(D|M)$ is the code length for encoding the transformed 3DGS attributes.
+In principle, increasing transform complexity can reduce $L(D|M)$, but simultaneously increases $L(M)$, potentially canceling out the compression gain.
+Optimizing the full MDL objective for TTC is highly challenging, because it requires non-trivial discrete–continuous optimization over the transform’s parameters and architecture.
+MDL optimization problem is only well posed for a given the model class.
+As the attainable trade-off between $L(M)$ and $L(D|M)$ is with respect to the chosen transform family, search-based solvers such as NAS (Elsken et al., 2019) can only explore what this search space permits.
+Thus, a more important and unanswered question is how to design an appropriate transform architecture for scene-specific 3DGS compression.
+Once such an architecture defines a suitable transform family as the search space, future work can readily apply existing NAS to search for MDL-efficient designs.
+In this work, we focus on this prerequisite step and propose an MDL-aware transform architecture that is parameter- and computation-efficient.
+To instantiate this MDL-aware, parameter- and computation-efficient transform design within TTC, we propose sparsity-guided hierarchical transform coding (SHTC).
+SHTC avoids computationally prohibitive repeated spatial KNN graph construction by restricting transforms to the channel domain, and uses a lightweight two-layer hierarchy to keep parameter and computational overhead low.
+In Layer 1, we apply the Karhunen-Loève Transform (KLT) to decorrelate channels and compact energy, yielding coefficients with highly uneven energy distribution.
+Coding only the principal coefficients reduces rate but introduces truncation error, whereas coding all coefficients is rate-inefficient.
+To balance this trade-off, Layer 2 serves as a refinement layer and uses a neural transform to compress the KLT residual, compensating for truncation-induced information loss with modest additional rate.
+Since the residual is typically compressible with many near-zero entries, we inject this sparsity prior as an inductive bias into the refinement layer, enabling effective residual coding with very few parameters.
+Concretely, we adopt a compressed-sensing-inspired refinement that encodes the residual with a small set of learned linear measurements and reconstructs it using a lightweight deep-unfolding decoder (Gregor and LeCun, 2010; Zhang et al., 2020; Zhang and Ghanem, 2018).
+Overall, we introduce TTC, which enables joint optimization of the 3DGS representation, the entropy model, and the transforms.
+Within this paradigm, we propose SHTC, a parameter- and computation-efficient transform architecture that achieves substantial R–D gains with minimal decoding-time overhead and moderate training-time cost; moreover, Fig.
+2 shows that it is empirically Pareto-optimal in the BD-rate–decoding-time trade-off compared with current state-of-the-art 3DGS compressors.
 In addition, our parameter-efficient design strategy, which combines a sparsity prior with deep unfolding, may offer a potential blueprint for developing low-complexity neural image and video codecs.
+Figure 2: Trade-off between decoding time and BD-rate (lower is better), with marker size indicating training time for HAC, ContextGS, CAT-3DGS, HAC++, and Ours.
+Figure 3: Illustration of (a) Scaffold-GS and (b) an anchor-based compression pipeline, using HAC as an example.
+2 Preliminary and Related Work
+In this work, to achieve the best compression performance, SHTC is integrated into anchor-based frameworks to enable the joint optimization of 3DGS, the entropy model, and the transforms.
+Since these anchor-based frameworks are built upon Scaffold-GS (Lu et al., 2024), we first recap the Scaffold-GS representation and then review how recent anchor-based methods further compress it.
+Due to space constraints, a more comprehensive introduction for the unstructured compression methods is deferred to Appendix G.1; a dedicated review of post-training transform coding in existing 3DGS compression approaches is presented in Appendix G.2.
+Scaffold-GS introduces a hierarchical structure during training: a sparse set of anchors serve as reference points from which a dense set of Gaussians are generated.
+As shown on the left side of Fig.
+3, each anchor is parameterized by the position $\mathbf{p}$, latent feature $\mathbf{f}$, scaling factor $\mathbf{l}$, and $K$ learnable offsets $\{\mathbf{O}_{i}\}_{i=1}^{K}$.
+The scaling factor $\mathbf{l}$ together with the offsets $\{\mathbf{O}_{i}\}_{i=1}^{K}$ determine the spatial distribution of the generated Gaussians.
+For each generated Gaussian, the view-dependent color, opacity, and covariance (parameterized by quaternion + scale) are predicted from the anchor’s latent feature through lightweight MLPs.
+Although Scaffold-GS is more compact than vanilla 3DGS, it primarily performs representation compaction.
+Anchor parameters and attributes are still stored as raw 32-bit floats, without quantization or entropy coding, which fundamentally limits the achievable bitrate reduction.
+SOGS reduces the dimensionality of anchor features (Zhang et al., 2025), but it still falls into the category of compaction rather than learned compression, and the gains remain limited.
+To overcome these limitations, recent anchor-based methods introduce context-based conditional entropy models to quantize and entropy-code anchor attributes, substantially improving compression ratios.
+HAC is an early representative baseline in this line of work.
+As illustrated in the right panel of Fig.
+3, HAC employs multi-scale hash grids as a hyperprior (Ballé et al., 2018).
+Given an anchor position, it queries the hash grids to obtain contextual features, which are then mapped by an MLP to the quantization step size $q_{s}$ and the Gaussian likelihood parameters $(\boldsymbol{\mu},\boldsymbol{\sigma})$.
+Building on HAC, subsequent methods improve the rate–distortion performance by adopting more expressive entropy models, such as channel-wise autoregressive model (ChARM) (Minnen and Singh, 2020) in HAC++ (Chen et al., 2025c) and CAT-3DGS (Zhan et al., 2025), as well as spatial autoregressive model (SARM) in ContextGS (Wang et al., 2024b) and HEMGS (Liu et al., 2024a).
+These methods represent the current state-of-the-art compression performance for 3DGS compression.
 (a) HAC
@@
 Figure 4: Comparison of inter-channel correlation (visualized as absolute Pearson correlation matrices) and energy compaction in the anchor features of HAC/HAC++ and their KLT coefficients, using the ‘playroom’ scene as an example.
-2 Related Work
-Following the taxonomy of [ali2025compression], we group existing 3DGS compression methods into (i) unstructured approaches, which employ model compression techniques such as pruning [ali2024elmgs, ali2024trimming, fan2024lightgaussian, niemeyer2024radsplat, lee2024compact, hanson2025pup] and quantization [girish2024eagles, navaneet2024compgs, wang2024end] for compression, and (ii) structured approaches [morgenstern2024compact, lee2025compression, lu2024scaffold], which exploit the spatial relationship among Gaussian primitives for compression.
-Because anchor-based designs are especially prevalent within structured approaches and our method builds on this architecture, we dedicate this section to anchor-based methods.
-Due to space limitations, a more detailed literature review is provided in the supplementary material.
-Among anchor-based 3DGS compression methods, Scaffold-GS [lu2024scaffold] is the representative baseline.
-Scaffold-GS introduces a hierarchical structure during training: a sparse set of anchors serve as reference points from which a dense set of Gaussians are generated.
-As shown on the left side of Fig.
-1, each anchor is parameterized by the position $\mathbf{p}$, latent feature $\mathbf{f}$, scaling factor $\mathbf{l}$, and $K$ learnable offsets $\{\mathbf{O}_{i}\}_{i=1}^{K}$.
-The scaling factor $\mathbf{l}$ together with the offsets determine the spatial distribution of the generated Gaussians.
-For each generated Gaussian, the view-dependent color, opacity, and covariance (parameterized by quaternion + scale) are predicted from the anchor’s latent feature through lightweight MLPs.
-Although Scaffold-GS is more compact than vanilla 3DGS, it is essentially a 3DGS compaction scheme: anchors and their attributes are still stored as raw 32-bit floats without quantization or entropy coding, which inherently limits the achievable compression ratio.
-SOGS reduces the dimensionality of anchor features; however, this remains 3DGS compaction rather than true compression, so the gains are still limited [zhang2025sogs].
-To address these limitations, recent work introduces context-based conditional entropy models that reduce statistical redundancy across anchor attributes, achieving significantly higher compression ratios.
-Among these methods, HAC was introduced early and serves as a foundational baseline.
-As illustrated in the right panel of Fig.
-1, HAC uses multi-scale hash grids as a hyperprior [balle2018variational].
-Given an anchor’s position, it queries these hash grids to obtain contextual features, which an MLP maps to the quantization step size $q_{s}$, the Gaussian likelihood parameters $(\boldsymbol{\mu},\boldsymbol{\sigma})$.
-HAC++ [chen2025hac++] enhances this design by (i) splitting each anchor feature vector into non-overlapping subgroups and encoding them sequentially with a channel-wise autoregressive model (ChARM) [minnen2020channel], and (ii) replacing a single Gaussian with a Gaussian mixture model to model more complex densities.
-CAT-3DGS [zhan2025catdgs] also adopts ChARM but substitutes hash-grid hyperpriors with triplanes.
-Both ContextGS [wang2024contextgs] and HEMGS [liu2024hemgs] follow a spatially sequential scheme: they use previously decoded anchors as conditioning context in a spatial autoregressive model (SARM) to encode the next anchors.
-These approaches represent the current state-of-the-art compression performance for 3DGS compression.
-3 Research Motivation and Challenges
-In this section, we analyze the importance and challenges of implementing end-to-end optimized transform coding in 3DGS compression.
+3 Method
+This section outlines the motivation and challenges for transitioning from anchor-based compression and PTC to our proposed TTC paradigm. We then present our main technical contribution, SHTC.
 3.1 Motivation
 Decorrelation
-Anchor-based compression methods, built on Scaffold-GS, achieve high compression ratios by quantizing and entropy-coding anchor attributes.
-Let $\mathbf{x}$ denote an anchor-attribute vector.
-With entropy coding, the ideal rate equals the source entropy $H_{q}=\mathbb{E}_{\mathbf{x}\sim q(\mathbf{x})}[-\log q(\mathbf{x})]$, but the ground truth distribution $q(\mathbf{x})$ is unknown.
-In practice, one can approximate the unknown distribution $q(\cdot)$ with a parametric model $p(\cdot)$ (often modeled as Gaussian), so the actual rate is the cross-entropy
-$R=\mathbb{E}_{\mathbf{x}\sim q(\mathbf{x})}[-\log p(\mathbf{x})]=H_{q}+D_{KL}(q||p).$ (1)
-where $D_{KL}(q||p)$ measures the extra bits caused by the mismatch between $q(\cdot)$ and $p(\cdot)$.
-Since anchor attributes are correlated, anchor-based methods use autoregressive models to exploit these correlations and reduce $D_{KL}$, thereby tightening the cross-entropy rate.
-For intra-vector correlation, $\mathbf{x}$ can be partitioned into groups of size $g$ and coded sequentially with the probability model $p(\mathbf{x}_{i:i+g}|\mathbf{x}_{<i})$ (e.g., ChARM as used in HAC++ [chen2025hac++] and CAT-3DGS [zhan2025catdgs]); for spatial correlation, anchors are coded causally with the probability model $p(\mathbf{x}^{(i)}|\{\mathbf{x}^{(j)}\}_{j\in\mathcal{C}_{i}})$ where $\mathcal{C}_{i}$ denotes the causal neighbors of anchor-$i$ (SARM as in ContextGS [wang2024contextgs] and HEMGS [liu2024hemgs]).
-While more complex entropy models can improve compression by exploiting data correlations during entropy coding, they also increase decoding latency.
-In 3DGS’s “compress once, decompress many times” pattern, decoding latency on the client side is critical; complex entropy models slow down decoding and degrade user experience.
-Moreover, even with more complex entropy models, fully capturing these dependencies remains difficult, leading to sub-optimal compression performance.
-These limitations arise from the absence of end-to-end optimized analysis–synthesis transforms.
-This motivates introducing such transforms into the compression pipeline: by decorrelating the signal before entropy coding, redundancy is reduced and much simpler, more efficient entropy models can be used.
-As shown in the first row of Fig.
-2, HAC and HAC++ anchor features exhibit strong inter-channel correlations.
-Applying a channel-wise KLT largely removes these correlations, yielding decorrelated coefficients and thereby paving the way for the use of simpler entropy models.
+Let $\mathbf{x}$ denote the anchor-attribute vector to be compressed.
+Under ideal entropy coding, the minimum achievable rate equals the source entropy $H_{q}=\mathbb{E}_{\mathbf{x}\sim q(\mathbf{x})}[-\log q(\mathbf{x})]$, but the true distribution $q(\mathbf{x})$ is typically unknown.
+In practice, one can approximate the unknown distribution $q(\cdot)$ with a parametric model $p(\cdot)$ (often modeled as Gaussian), leading to the cross-entropy rate
+$R=\mathbb{E}_{\mathbf{x}\sim q(\mathbf{x})}[-\log p(\mathbf{x})]=H_{q}+D_{KL}(q||p)$ (1)
+where $D_{KL}(q||p)$ measures the extra bits due to model mismatch.
+Anchor attributes exhibit strong statistical dependencies; assuming independence across dimensions/channels therefore increases mismatch and inflates the rate.
+Existing anchor-based methods mitigate this issue by using correlation-aware entropy models (e.g., ChARM or SARM) to better capture dependencies and reduce $D_{KL}$.
+However, complex entropy models typically incur higher decoding cost (often due to sequential/context-dependent decoding), which is undesirable in 3DGS’s “compress once, decompress many times” setting where client-side decoding latency is critical.
+Moreover, fully modeling complex dependencies remains challenging even with sophisticated entropy models, limiting compression gains.
+A complementary approach is to introduce an analysis transform and decorrelate $\mathbf{x}$ before entropy coding.
+By reducing redundancy in the representation, we can rely on simpler and faster entropy models without sacrificing rate performance.
+As illustrated in the first row of Fig.
+4, HAC and HAC++ features show pronounced inter-channel correlations; a more complex entropy model (as in HAC++) models these dependencies but does not remove them from the signal.
+In contrast, applying a channel-wise KLT substantially suppresses inter-channel correlations, producing near-decorrelated coefficients and paving the way for the use of simpler entropy models.
 Energy compaction
@@
 This effect is evident in the second row of Fig.
-2, where KLT compacts most of the energy into the principal coefficients.
-Applying an analysis transform before entropy coding decorrelates the data and compacts its energy, which benefits lossy compression, while a paired synthesis transform reconstructs the anchor attributes for rendering.
-When the transforms and entropy models are jointly optimized, this transform coding paradigm typically achieves superior R–D performance, as demonstrated in neural image compression [balle2016end, balle2018variational, minnen2018joint, cheng2020learned, he2022elic, zou2022devil, zhu2021transformer, liu2023learned, ali2024towards].
-Motivated by these successes, extending this paradigm to 3DGS compression is a natural choice.
-3.2 Challenges
-Despite the promise of end-to-end optimized transform coding, to the best of our knowledge, no existing work has applied such transforms to 3DGS compression.
-Extending end-to-end optimized transform coding from neural image compression to 3DGS is, however, non-trivial due to two domain-specific challenges.
-The first challenge stems from the irregular, unordered nature of 3DGS anchor points, which lack a natural neighborhood structure.
-Although neighborhoods can be defined via KNN, repeated neighbor searches in dense point clouds are computationally expensive.
-This issue is exacerbated by the fact that anchor positions are optimization variables that evolve during training, so the KNN graph must be rebuilt throughout optimization rather than constructed once, significantly increasing the overall training cost.
-The second challenge arises from a fundamental difference between 3DGS compression and neural image/video codecs.
-In standard neural codecs, the analysis–synthesis transforms are trained once, shared across inputs, and omitted from the bitstream; consequently, large-capacity transforms can be used to pursue better R-D trade-offs.
-By contrast, 3DGS relies on scene-specific transforms whose parameters must be transmitted, imposing a strict parameter budget.
-Under this constraint, strong R–D performance hinges on architectural efficiency rather than sheer model size.
-Historically, early neural image codecs with relatively simple transforms [balle2016end, balle2018variational] failed to surpass traditional codecs, whereas later models with substantially more complex transforms did [cheng2020learned].
-This evolution highlights the difficulty of designing neural transforms that are both parameter-efficient and capable of strong R-D performance.
-These two challenges call for careful architectural design to balance R-D gains against additional computational and parameter costs.
-To address them, we propose SHTC, a transform-coding framework for 3DGS that is both computationally and parameter efficient, achieving strong R-D performance under a tight parameter budget.
-Technical details of SHTC are provided in Sec.
-4.
-4 SHTC: Sparsity-guided Hierarchical Transform Coding
-[…]
-4.1 Layer 1: Base Layer
+4, where KLT compacts most of the energy into the principal coefficients.
+PTC vs. TTC: decoupled vs. co-adaptive optimization.
+Applying an analysis transform before entropy coding can decorrelate the signal and concentrate its energy into fewer coefficients, which improves the R-D efficiency of lossy compression.
+A paired synthesis transform then reconstructs the anchor attributes for rendering.
+A few prior works have also explored analysis-ynthesis transforms (Morgenstern et al., 2024; Lee et al., 2025b), but they typically follow the PTC paradigm (see Fig.
+1): applying a fixed, pre-defined transform codec to a frozen 3DGS model in the post-training stage (details in Sec.
+G.2).
+This two-stage, decoupled design prevents mutual adaptation between the transform and the underlying 3DGS representation, often leading to sub-optimal R-D performance.
+In contrast, we advocate the TTC paradigm, where the transforms and entropy model are trained jointly with the 3DGS representation, allowing them to co-adapt and yielding a more compressible representation without sacrificing rendering quality.
+3.2 Challenges and Design Considerations
+TTC is an open-ended paradigm with many candidate transforms (e.g., DCT/DWT, generic MLPs, and graph-based spatial transforms).
+A natural question then arises: what kinds of transforms are appropriate?
+To answer this question, we need to understand two domain-specific challenges in 3DGS compression.
+The first challenge stems from the irregular, unordered nature of anchor/Gaussian primitives, which lack a natural neighborhood structure.
+Although neighborhoods can be defined via K-Nearest Neighbors (KNN), neighbor searches in dense point clouds are expensive.
+Moreover, anchor/Gaussian positions are optimization variables and thus change during training, precluding a one-time KNN construction and requiring repeated graph rebuilding, which substantially increases training cost.
+This makes graph-based spatial transforms prohibitively expensive to integrate into end-to-end 3DGS training.
+The second challenge arises from a fundamental difference between 3DGS compression and neural image/video codecs (Ballé et al., 2018; Minnen et al., 2018; Cheng et al., 2020; He et al., 2022; Li et al., 2024).
+Standard neural codecs employ a single analysis-synthesis transform shared across inputs, and its parameters are not included in the bitstream; this allows the use of parameter-heavy transforms to improve the R–D performance.
+In contrast, under the TTC paradigm, 3DGS relies on scene-specific transforms whose parameters must be transmitted, imposing a strict parameter budget.
+Consequently, R-D performance hinges on architectural efficiency and cannot be improved simply by increasing the number of model parameters.
+Under a tight parameter budget, purely linear transforms often lack sufficient expressive power, while purely MLP-based transforms may be ineffective without suitable inductive bias.
+In light of these constraints, we propose SHTC.
+SHTC circumvents the first challenge by restricting the transform to be channel-wise, i.e., without spatial interactions, making it computation-friendly.
+To remain effective under a low parameter budget, SHTC adopts a hierarchical architecture, with interpretable components.
+Details are provided in Sec.
+3.3.
+3.3 SHTC: Sparsity-guided Hierarchical Transform Coding
 Transform coding targets decorrelation and energy compaction; among orthonormal linear transforms, KLT is optimal for both.
-Under a tight parameter budget, we also find that an MLP-based transform lacks the capacity to surpass the KLT.
+Under a tight parameter budget, we also observe that an MLP-based transform fails to outperform the KLT (see Appendix C).
 Overall, the KLT offers an excellent cost–performance trade-off, achieving strong R-D performance with minimal compute and memory.
-We therefore adopt the KLT in Layer 1.
-Let $\mathbf{f}\in\mathbb{R}^{N_{f}}$ denote the anchor feature, with global mean $\mathbf{m}\in\mathbb{R}^{N_{f}}$ and KLT basis $\mathbf{V}\in\mathbb{R}^{N_{f}\times N_{f}}$.
-The analysis transform is $\boldsymbol{\theta}=g_{a}(\mathbf{f})=\mathbf{V}^{\top}(\mathbf{f}-\mathbf{m})$.
+Therefore, we use the KLT as the foundation for constructing our transform.
 The KLT coefficients exhibit an uneven energy distribution.
-For the anchor feature, we retain only the top-$M$ principal coefficients $\boldsymbol{\theta}_{p}=\boldsymbol{\theta}_{1:M}$, which are quantized and entropy-coded.
-The base layer reconstruction is $\hat{\mathbf{f}}_{base}=\mathbf{V}_{:,1:M}\hat{\boldsymbol{\theta}}_{p}+\mathbf{m}$.
-4.2 Layer 2: Refinement Layer via Sparsity-Aware Residual Coding
-Discarding the remaining coefficients causes information loss.
+Discarding low-energy coefficients causes information loss.
 The artifacts introduced by this truncation become more pronounced at high rates, leading to a quality bottleneck.
 However, coding all coefficients is rate-inefficient.
-To mitigate this information loss while keeping the rate overhead modest, we introduce a neural refinement layer.
-This layer transmits the KLT residual to compensate for truncation.
-Because most of the signal energy is captured by the leading KLT coefficients, the residual is dominated by low-magnitude components and can be well approximated as (approximately) sparse.
-For sparse or compressible signals, compressed sensing theory shows that the full signal can be reliably reconstructed from a small number of linear measurements by solving a sparsity-regularized optimization problem [donoho2006compressed].
-This insight motivates us to use a principled, sparsity-aware design for the refinement transforms.
-Instead of relying on a generic multilayer perceptron (MLP), we adopt a parameter-efficient transform inspired by compressed sensing and deep unfolding techniques [zhang2020deep, zhang2018ista], enabling flexible compression of the residual with low parameter cost.
+To balance these two extremes, we propose a hierarchical framework integrating KLT with a neural refinement layer.
+  • Layer 1: We use the KLT to construct a base layer for coarse reconstruction.
+Let $\mathbf{f}\in\mathbb{R}^{N_{f}}$ denote the anchor feature, with global mean $\mathbf{m}\in\mathbb{R}^{N_{f}}$ and KLT basis $\mathbf{V}\in\mathbb{R}^{N_{f}\times N_{f}}$.
+The base-layer analysis transform is given by $\boldsymbol{\theta}=g_{a}(\mathbf{f})=\mathbf{V}^{\top}(\mathbf{f}-\mathbf{m})$.
+For the anchor feature, we retain only the top-$M$ principal coefficients $\boldsymbol{\theta}_{p}=\boldsymbol{\theta}_{1:M}$, which are quantized and entropy-coded.
+The base-layer synthesis transform is given by $\hat{\mathbf{f}}_{base}=g_{s}(\hat{\boldsymbol{\theta}}_{p})=\mathbf{V}_{:,1:M}\hat{\boldsymbol{\theta}}_{p}+\mathbf{m}$.
+  • Layer 2: To compensate for the truncation error of the base layer without incurring significant rate overhead, we introduce a neural refinement layer to transmit the KLT residual.
+Because most of the signal energy is captured by the leading KLT coefficients, the residual is is typically low-magnitude and approximately sparse/compressible.
+This observation raises a natural question: can we leverage this sparsity prior to introduce an inductive bias, enabling a very small refinement network to perform effective residual coding with strong R–D performance?
+We answer this question affirmatively by drawing on classical compressed sensing (CS), which suggests that a high-dimensional sparse (or compressible) signal can often be reconstructed from a small number of linear measurements via sparsity-regularized recovery (Donoho, 2006).
+Motivated by this insight, we design the refinement transforms in a sparsity-aware, CS-inspired manner.
+Specifically, the refinement-layer analysis transform produces a compact set of learned linear measurements of the residual, while the refinement-layer synthesis transform reconstructs the residual from the decoded measurements using an ISTA-style deep unfolding approach under a sparsity prior (Zhang and Ghanem, 2018; Zhang et al., 2020).
+A detailed description of these two transforms is provided next.
 3.3.1 Analysis Transform: Learned Linear Measurements of Residuals
@@
 The measurement vector $\mathbf{y}$ is then quantized to $\hat{\mathbf{y}}$ and entropy coded.
-4.2.2 Synthesis Transform: Sparsity-Regularized Residual Reconstruction
+3.3.2 Synthesis Transform: Residual Reconstruction
 Given the decoded quantized measurements $\hat{\mathbf{y}}$, the synthesis transform $h_{s}(\cdot)$ reconstructs an estimate $\hat{\mathbf{r}}$ of the residual:
@@
 This inverse problem can be addressed using a variety of well-established optimization algorithms.
-A classical choice is the Iterative Shrinkage-Thresholding Algorithm (ISTA) [zibulevsky2010l1].
+A classical choice is the Iterative Shrinkage-Thresholding Algorithm (ISTA) (Zibulevsky and Elad, 2010).
 At the $k$-th iteration, ISTA updates the current estimate by performing a gradient descent step on the quadratic data fidelity term, followed by a soft-thresholding operation to handle the non-smooth $\ell_{1}$ regularization term.
@@
 Directly applying ISTA, however, can be computationally expensive, as it typically requires many iterations and manual tuning of the hyperparameters $\tau$ and $\eta$.
-To address this, we adopt a deep unfolding approach [gregor2010learning, zhang2018ista, zhang2020deep] that maps a large number of ISTA iterations to a small number of learnable network layers, significantly reducing inference time while preserving interpretability and the sparsity prior.
+To address this, we adopt a deep unfolding approach (Gregor and LeCun, 2010; Zhang and Ghanem, 2018; Zhang et al., 2020) that maps a large number of ISTA iterations to a small number of learnable network layers, significantly reducing inference time while preserving interpretability and the sparsity prior.
 Each layer corresponds to one unfolded ISTA iteration, with layer-wise learnable step sizes and thresholds.
@@
 Finally, the predicted residual $\hat{\mathbf{r}}$ is combined with the base layer reconstruction to form the final anchor feature for rendering.
+Figure 5: Illustrative system pipeline integrating SHTC into HAC.
+Anchor coordinates are encoded by MPEG-GPCC, while SHTC latents and other attributes are quantized and entropy-coded using a hash-assisted context model that predicts probability parameters and quantization step sizes conditioned on $\mathbf{p}$.
 3.4 Implementation details
-There are four types of objects that we need to compress:
-  • Anchor positions: following HAC++ [chen2025hac++], we compress anchor coordinates using MPEG-GPCC [graziosi2020overview].
-  • Features $\mathbf{f}\in\mathbb{R}^{50}$: we compress feature vectors with the full two-layer SHTC.
-The base layer encodes $\mathbf{f}$ into a low-dimensional representation $\boldsymbol{\theta}_{p}$ with $M=15$.
-The refinement layer then encodes the KLT residual $\mathbf{r}$ into a measurement vector $\mathbf{y}$ of dimension $N_{s}=15$.
-  • Scalings $\mathbf{l}\in\mathbb{R}^{6}$: for the scaling vectors, we apply only the first SHTC layer (KLT) without dimensionality reduction.
-Because $\mathbf{l}$ is low-dimensional and the reconstruction quality is highly sensitive to it, truncating KLT coefficients in the transform domain severely degrades performance.
-Instead, for the KLT coefficients $\mathbf{l}_{t}\in\mathbb{R}^{6}$ we adopt a monotonically increasing, channel-wise quantization schedule:
-    $q_{s}^{(i)}=q_{s}\exp(\alpha i),\quad i\in\{0,\ldots,5\},$ (8)
-    where $\alpha$ is a learnable scalar and $q_{s}$ is the per-anchor base quantization step predicted by HAC.
-This quantization schedule acts as a remedial measure when dimensionality reduction cannot be applied: it assigns larger steps to lower-energy components, improving the rate–distortion trade-off while avoiding truncation loss.
-  • Offsets $\{\mathbf{O}_{i}\}_{i=1}^{K}$: for offset vectors, we retain the original treatment in HAC.
-Due to offset masking, offset vectors have variable dimensionality across anchors, which makes a fixed transform impractical.
-As described in Sec.
-2 and illustrated in Fig.
-1, HAC uses multi-scale hash grids to parameterize its entropy model.
-The original HAC queries multi-scale hash grids to predict the quantization step $q_{s}$ and Gaussian likelihood parameters $(\boldsymbol{\mu},\boldsymbol{\sigma})$ for three types of anchor attributes: $\mathbf{f}$, $\mathbf{l}$, and $\{\mathbf{O}_{i}\}_{i=1}^{K}$.
-After integrating SHTC into HAC, we slightly modify the entropy coding process.
-It still queries multi-scale hash grids, but now predicts $q_{s}$ and $(\boldsymbol{\mu},\boldsymbol{\sigma})$ for the transformed vectors $\boldsymbol{\theta}_{p}$, $\mathbf{y}$, and $\mathbf{l}_{t}$, as well as for the offset vectors $\{\mathbf{O}_{i}\}_{i=1}^{K}$.
-The remaining quantization and coding procedures are the same as in the original HAC.
-Inspired by color-space transforms in traditional codecs, we replace the original pixel-wise $\ell_{1}$ loss between the rendered image $\hat{I}$ and ground truth $I$ with a YCbCr-based loss, $\mathcal{L}_{\mathrm{YCbCr}}$, which computes reconstruction errors in the YCbCr color space.
-Because the human visual system is more sensitive to luminance than to chrominance variations, we assign a higher weight to the Y component and add a sharpness term on Y together with a smoothness term on Cb/Cr to preserve structure while suppressing color noise.
-This loss regularizes the compressed 3DGS representation, yielding fewer luminance errors and lower perceptual distortion at a given bitrate and thereby modestly improving R–D performance.
-The full definition of $\mathcal{L}_{\mathrm{YCbCr}}$ is provided in the supplementary material.
-The overall training loss is defined as
-$\mathcal{L}=\mathcal{L}_{d}+\lambda\mathcal{L}_{\mathrm{b-rate}}+\lambda_{o}\mathcal{L}_{\mathrm{reg}}+\lambda_{e}\ell_{1}(\mathbf{r},\hat{\mathbf{r}})+\lambda_{r}\mathcal{L}_{\mathrm{r-rate}}$ (9)
-[…]
-5 Experimental results
-5.1 Experimental setup
-Dataset
-We primarily evaluate the R-D performance on three commonly used large-scale real-scene datasets: DeepBlending [hedman2018deep], Mip-NeRF360 [barron2022mip], and Tanks&Temples [knapitsch2017tanks].
-In particular, we assess all nine scenes from the Mip-NeRF360 dataset [barron2022mip].
-Evaluations on additional datasets are provided in the supplementary material.
-These diverse datasets provide a comprehensive evaluation of the proposed SHTC framework.
-Distortion metrics
-We evaluate rendering distortion using three metrics: PSNR, SSIM [wang2004image], and LPIPS [zhang2018unreasonable].
-Rate metrics
-We adopt the BD-rate metric [bjontegaard2001calculation] to quantify memory savings by comparing the R–D curve of a compression method against that of a baseline.
-Specifically, BD-rate computes the average bitrate difference between the two R–D curves over a common distortion range.
-A negative BD-rate indicates that, on average, the method achieves a lower bitrate (measured in memory size) than the baseline at comparable distortion levels, with the absolute value (expressed as a percentage) quantifying the average bitrate reduction.
-Training details
-We follow the official training settings of HAC [chen2024hac] and HAC++ [chen2025hac++], and train each scene for 30k iterations.
-The Lagrange multiplier $\lambda$ is set to $\{0.002,0.004,0.008,0.015\}$ to obtain R–D curves that span a wide range of bitrates and distortions for a more comprehensive comparison.
-The gaps between successive $\lambda$ values are chosen to increase because the effect of $\lambda$ on the achieved bitrate is highly non-linear: especially in the low-bitrate regime, larger changes in $\lambda$ are required to produce clearly separated bitrates and thus more evenly distributed points on the R–D curve.
-We set $\lambda_{e}=0.03$ and $\lambda_{r}=\max(\lambda/4,0.001)$.
-All experiments are conducted on a server equipped with two RTX 4090 GPUs.
-5.2 Comparison with Anchor-based Methods
-5.2.1 Evaluating R-D performance
+We integrate the proposed SHTC into the HAC framework to build our 3DGS compressor.
+In Fig.
+5, we illustrate how the SHTC analysis transform is integrated into the HAC pipeline before quantization and entropy coding.
+The synthesis transform mirrors the analysis transform and is omitted for brevity.
+Following HAC++ (Chen et al., 2025c), anchor coordinates $\mathbf{p}$ are compressed using MPEG-GPCC (Graziosi et al., 2020), yielding a separate coordinate bitstream.
+For the offset vectors $\{\mathbf{O}_{i}\}_{i=1}^{K}$, we keep the original HAC design.
+Due to offset masking, the number of valid offsets varies across anchors, making a fixed-dimensional transform impractical; therefore $\{\mathbf{O}_{i}\}_{i=1}^{K}$ are directly quantized and entropy coded as in HAC.
+For the scaling vector $\mathbf{l}$, we use only the base layer of SHTC (without dimensionality reduction) to produce $\mathbf{l}_{t}\in\mathbb{R}^{6}$, and then apply channel-wise step-size modulation during quantization following Eq.(8) with a dimension-adapted formulation, so that principal coefficients use smaller step sizes than less important ones; the resulting symbols are entropy-coded by HAC.
+For the feature vector $\mathbf{f}\in\mathbb{R}^{50}$, the full SHTC produces two latent representations, $\boldsymbol{\theta}_{p}$ and $\mathbf{y}$, with dimensions $M=15$ and $N_{s}=15$, respectively; both are quantized and entropy coded by HAC.
+As shown in Fig.
+5, the probability parameters (e.g., mean and scale) and the quantization step $q_{s}$ for $\boldsymbol{\theta}_{p}$, $\mathbf{y}$, $\mathbf{l}_{t}$ and $\{\mathbf{O}_{i}\}_{i=1}^{K}$ are predicted from the multi-scale hash grids conditioned on the query position $\mathbf{p}$, and are used by the quantizer and the entropy coder.
+We follow an HAC-style training objective and, in implementation, compute the pixel-wise reconstruction error in YCbCr space rather than RGB.
+Appendix E reports the full loss formulation and hyperparameter settings; the YCbCr distortion term is detailed in Appendix E.1.
 (a) Mip-NeRF360
@@
 Figure 6: Comparison of our method with existing anchor-based 3DGS compression methods.
-Fig.
-3 shows the R-D curves of several anchor-based 3DGS compression methods.
+4 Experimental Results
+4.1 Experimental Setup
+[…]
+Dataset
+We evaluate the R-D performance on five commonly used large-scale real-scene datasets: DeepBlending (Hedman et al., 2018), Mip-NeRF360 (Barron et al., 2022), Tanks&Temples (Knapitsch et al., 2017), Synthetic-NeRF (Mildenhall et al., 2021) and BungeeNeRF (Xiangli et al., 2022).
+In particular, we assess all nine scenes from the Mip-NeRF360 dataset (Barron et al., 2022).
+Evaluations on additional datasets are provided in the supplementary material.
+These diverse datasets provide a comprehensive evaluation of the proposed SHTC framework.
+Distortion metrics
+We evaluate rendering distortion using three metrics: PSNR, SSIM (Wang et al., 2004), and LPIPS (Zhang et al., 2018).
+Rate metrics
+Following prior work, we use the storage cost (measured as memory size in MB) as the rate metric.
+We adopt the BD-rate (Bjontegaard, 2001) to quantify rate savings by comparing the R-D curve of a method against a baseline.
+Specifically, BD-rate reports the average rate difference (in percentage) between two R-D curves over their overlapping distortion range.
+A negative BD-rate indicates that the method requires less storage than the baseline at comparable distortion levels, and the magnitude reflects the average percentage of rate reduction.
+4.2 Comparison with Anchor-based Methods
+4.2.1 Evaluating R-D performance
+Figure 6 shows the R-D curves of several anchor-based 3DGS compression methods.
 Across all three datasets, the curve corresponding to the proposed SHTC, labeled “Ours SHTC (Full)” in the legend, lies toward the upper-left region, indicating higher PSNR at a given size or, equivalently, smaller size at a given distortion.
-For PCGS [chen2025pcgs], we train the model with $\lambda\in\{0.004,0.008,0.015,0.025\}$ to span a wide distortion range, so that its R-D points exhibit sufficient PSNR variation and a modest amount of overlap with the other methods, making the differences in the bitrate required to reach a given PSNR clearly visible.
-For HEMGS [liu2024hemgs], which is not open-sourced, we plot the R-D points reported in their paper and conduct an approximate comparison.
-Using only half as many training iterations (30k vs 60k), our method performs on par with or better than HEMGS on Mip-NeRF360 and DeepBlending and appears competitive on Tanks&Temples.
-Tab.
-1 reports the corresponding BD-rate of SHTC relative to HAC++, CAT-3DGS, and ContextGS, quantifying the bitrate (memory footprint) savings at equal PSNR.
-On average over the three datasets, our method achieves BD-rate values of about -21.0%, -24.3%, and -43.6% relative to HAC++, CAT-3DGS, and ContextGS, respectively.
-Taken together, these results indicate that combining end-to-end optimized transforms with a simple entropy model is more effective than solely increasing entropy-model complexity in anchor-based 3DGS compression.
-This highlights the importance of incorporating transform coding and demonstrates the strong performance of the proposed SHTC framework.
-Dataset HAC++ CAT-3DGS (40k) ContextGS Mip-NeRF360 -20.81% -24.54% -49.36% Tank&Temples -22.55% -22.78% -39.17% DeepBlending -19.58% -25.53% -42.30%
-Table 1: The BD-rate of SHTC relative to three baselines: HAC++ [chen2025hac++], CAT-3DGS [zhan2025catdgs] and ContextGS [wang2024contextgs].
-Negative values indicate bitrate savings (better) compared with each baseline.
-5.2.2 Evaluating Computational Cost
-[…]
-ContextGS HAC++ Ours ContextGS HAC++ Ours ContextGS HAC++ Ours ContextGS HAC++ Ours
-Mip-NeRF360 106 128 146 32.73 8.48 5.90 33.32 13.86 8.86 -30.92% -56.60% -64.82%
-Tank&Temples 130 158 177 29.37 5.63 3.91 29.05 8.81 5.80 -26.41% -44.87% -56.06%
-DeepBlending 161 169 185 16.78 2.98 2.30 16.63 4.41 3.18 -41.09% -58.34% -64.56%
-Table 2: Evaluating inference efficiency and compression effectiveness: complex entropy modeling alone vs.
-SHTC with a simple entropy model.
-From left to right, we report four metrics: rendering FPS, encoding time, decoding time, and BD-rate relative to the HAC baseline.
-All results are obtained from models trained with $\lambda=0.004$.
-The best result for each metric is highlighted in bold.
-As shown in Tab.
-2, our method not only achieves greater BD-rate gains relative to the HAC baseline than HAC++ and ContextGS, but also reduces both encoding and decoding latency.
-This efficiency gain stems from a paradigm shift: instead of relying on a highly complex entropy model to capture dependencies among high-dimensional anchor features, SHTC uses learned transforms to explicitly decorrelate anchor attributes before entropy coding.
-By shifting the decorrelation burden from the entropy model to the transform, SHTC produces coefficients for which a simple and fast entropy model is sufficient, yielding better overall compression efficiency and lower latency than relying on a highly complex entropy model alone.
-In terms of training time, SHTC introduces a modest overhead.
-In our experiments, compared with HAC++, the training time per scene increases by at most about 8 minutes, while HAC++ itself already requires approximately 40–50 minutes to train.
-However, this additional cost is less critical in practice, because training is a one-time offline cost borne by 3DGS asset producers, whereas decoding is performed repeatedly on the user side: end users neither observe nor care about the training time, but they are directly sensitive to download and decoding latency.
-SHTC targets this more critical bottleneck by producing smaller bitstreams at comparable distortion and reducing decoding latency, making it particularly well suited for consumer-centric applications where fast access is essential.
-A detailed comparison of training times is provided in the supplementary material for reference.
-In terms of parameter overhead, the transforms in SHTC introduce 5,093 additional parameters over the HAC baseline.
+For HEMGS (Liu et al., 2024a), which is not open-sourced, we plot the R-D points reported in their paper and conduct an approximate comparison.
+Using only half the training iterations of HEMGS (30k vs.
+60k), our method performs on par with or better than HEMGS.
+The R-D curves tested on Synthetic-NeRF and BungeeNeRF are provided in Fig.
+7 in Appendix.
+Table 1 reports the corresponding BD-rate of our method (HAC + Ours SHTC) relative to HAC++, CAT-3DGS, ContextGS and HAC, quantifying the average memory footprint savings at equal PSNR.
+These results demonstrate the superiority of our proposed TTC paradigm over anchor-based compression, which relies solely on increasingly complex entropy models.
+Note that SHTC is not tailored to HAC and does not depend on it.
+When integrated into ContextGS, SHTC also yields consistent improvements over the vanilla ContextGS baseline.
+The R-D curves are shown in Fig.
+8 in Appendix, and ContextGS+SHTC achieves a BD-rate of $19.45\%$ relative to vanilla ContextGS.
+Table 1: The BD-rate of our method relative to four baselines: HAC++, CAT-3DGS, ContextGS and vanilla HAC. Negative values indicate bitrate savings (better) compared with each baseline.
+Dataset HAC++ CAT-3DGS (40k) ContextGS HAC Mip-NeRF360 -20.81% -24.54% -49.36% -64.82% Tank&Temples -22.55% -22.78% -39.17% -56.06% DeepBlending -19.58% -25.53% -42.30% -64.56% BungeeNeRF -10.04% -39.90% - -52.05% Synthetic-NeRF -13.45% - - -26.33%
+Table 2: Efficiency and rate–distortion comparison on Mip-NeRF360 ($\lambda=0.004$), evaluated on the same RTX 4090 server.
+HAC is reported as a reference anchor and is not included in the comparison.
+We compare Ours against ContextGS, CAT-3DGS, and HAC++ in terms of training time (in seconds), rendering FPS, encoding/decoding latency (in seconds), and BD-rate measured relative to the HAC baseline.
+Best results among the compared methods are highlighted in bold.
+Methods Training Rendering Encoding Decoding BD-rate Time FPS Time Time HAC 1949 131 4.20 10.05 0 ContextGS 3927 106 32.73 33.32 -30.92% CAT-3DGS 5725 - 79.82 87.95 -53.48% HAC++ 2735 128 8.48 13.86 -56.60% Ours 3109 146 5.85 8.84 -64.82%
+4.2.2 Evaluating Transform Overhead
+Computational overhead.
+We use vanilla HAC as the reference baseline and compare our method with state-of-the-art anchor-based codecs, including HAC++, ContextGS, and CAT-3DGS.
+We report BD-rate relative to HAC to quantify R-D gains, and we additionally measure encoding/decoding latency and training time to characterize the cost of these improvements.
+Results are summarized in Table 2 and visualized in Fig.
+2.
+Overall, our method achieves improved R–D performance while maintaining low coding latency.
+We attribute this advantage to the introduced transforms, which produce features that better match a lightweight, parallel-friendly entropy model, thereby avoiding expensive and complex entropy modeling.
+In terms of training cost, SHTC incurs substantially less overhead than CAT-3DGS and ContextGS, while being moderately slower than HAC++.
+Since many 3DGS deployments follow a “compress once, decompress many times” workflow, this training overhead is a one-time offline cost for asset producers and typically does not affect client-side user experience.
+In contrast, the memory footprint and decoding latency directly impact client-side user experience and are therefore the primary practical constraints.
+Consequently, the proposed SHTC method offers a favorable trade-off and lies on (or close to) the empirical Pareto frontier among the compared anchor-based methods as shown in Fig.
+2.
+Parameter overhead.
+The transforms in SHTC introduce 5,093 additional parameters over the HAC baseline.
 However, when SHTC is integrated into HAC, the entropy model queries the hash grids and feeds an MLP that predicts probability parameters for low-dimensional transform coefficients instead of high-dimensional anchor features.
@@
 In contrast, HAC++ increases the number of parameters by 45,400 relative to HAC.
-These results highlight the parameter efficiency of SHTC, which stems from a sparsity-aware transform design rather than simply adopting a larger black-box MLP.
+These results highlight the parameter efficiency of SHTC, which stems from the sparsity-aware transform design rather than simply adopting a larger black-box MLP.
 4.2.3 Visual Comparison
-Due to space limitations, visual comparisons are provided in the supplementary material.
-(a) Mip-NeRF360
-(b) Tank&Temples
-(c) DeepBlending
-Figure 4: Comparison of our method with several conceptually representative baslines.
+Due to space limitations, visual comparisons are provided in Appendix A.
 4.3 Comparison with Other Methods
-In addition to the comparisons with anchor-based compression methods using R–D curves and BD-rate, we also evaluate our method against a broad set of anchor-free and hybrid baselines [kerbl20233d, lu2024scaffold, lee2024compact, niedermayr2024compressed, girish2024eagles, fan2024lightgaussian, morgenstern2024compact, navaneet2024compgs, papantonakis2024reducing, wang2024end, liu2024compgs, chen2025fast, shin2025localityaware, tang2025neuralgs, hanson2025pup, xie2025sizegs, tian2025flexgaussian, lee2025optimized, tang2025feature, ma2025enhancing, morgenstern2024compact, lee2025compression, xie2024mesongs, yang2025hybridgs].
-Because many of these methods only report one or two operating points on the R-D curve, they do not support a meaningful BD-rate computation; for these methods, we therefore provide a numerical comparison in tabular form in the supplementary material.
-[…]
+In addition to the comparisons with anchor-based compression methods using R-D curves and BD-rate, we also evaluate our method against a broad set of anchor-free and hybrid baselines.
+Because many of these methods only report one or two operating points on the R-D curve, they do not support a meaningful BD-rate computation; for these methods, we therefore provide a numerical comparison in tabular form in Appendix D and Table 4 in Appendix.
+In addition, we provide a dedicated comparison with several conceptually representative baselines: the classic 3DGS (Kerbl et al., 2023) and Scaffold-GS (Lu et al., 2024), the current state-of-the-art (SOTA) anchor-free compression method OMG (NeurIPS 2025) (Lee et al., 2025a), the feed-forward compression method FCGS (Chen et al., 2025b), and four representative methods that apply transform coding in a post-training stage for 3DGS compression: SOG and MesonGS (ECCV 2024) (Morgenstern et al., 2024; Xie et al., 2024), CodecGS (ICCV 2025) (Lee et al., 2025b), HybridGS (ICML 2025) (Yang et al., 2025).
+Note that FCGS can also be seen as a method that applies transform coding in a post-training stage after 3DGS is fixed.
+As shown in Fig.
+13 in Appendix, our method consistently outperforms the above eight methods.
+These results clearly demonstrate that jointly optimizing the GS representation, entropy models, and the transforms in an end-to-end manner is substantially more effective than the ad hoc post-training transform paradigm where 3DGS and the transforms cannot be mutually adapted.
 4.4 Ablation Study
-As shown in Fig.
-3, using only the base layer of SHTC, denoted as ‘Ours SHTC (BL)’, achieves slightly better or comparable R–D performance compared to HAC++ at low and medium bitrates.
+As shown in Figure 6, using only the base layer of SHTC, denoted as ‘Ours SHTC (BL)’, achieves slightly better or comparable R–D performance compared to HAC++ at low and medium bitrates.
 However, the improvements are modest overall, and SHTC (BL) fails to consistently surpass HAC++ in the high-rate region.
@@
 At high bitrates, it also remains clearly superior in terms of R-D performance and does not suffer from the high-rate quality bottleneck observed with the base-layer-only variant.
-The supplementary material provides a more detailed quantitative ablation study and a further analysis of the choice of transforms, and also discusses the limitations of our approach and directions for future work.
+We conduct a thorough ablation study to isolate the contribution of each component; the main results are reported in Fig.
+10(a) and Table 3 in Appendix, with additional details deferred to Appendix B.1.
+In addition to component-wise effects, a natural question is whether residual transform coding is actually necessary for handling truncation.
+To answer this, we compare (i) our residual transform coding scheme that compensates truncation error against (ii) a direct baseline that transmits all KLT coefficients without truncation.
+This comparison is provided in Appendix B.2.
+Finally, since SHTC introduces a new transform design, one may wonder why we adopt this particular design instead of more standard alternatives (e.g., DCT, DWT, or a generic MLP).
+We therefore include a dedicated analysis of these transform choices in Appendix C.
+Limitations and future work are discussed in Appendix F.
 5 Conclusion
-In this paper, we introduced SHTC, to our knowledge, the first end-to-end optimized transform coding framework for 3DGS compression, which achieves substantial R–D gains and improved inference efficiency with only minimal parameter overhead.
-By leveraging learned analysis–synthesis transforms for decorrelation instead of relying solely on highly complex entropy models, SHTC improves compression efficiency and reduces decoding latency, making it well suited to consumer-facing scenarios where fast transmission and decoding are critical.
-SHTC has the potential to inspire future work on 3DGS compression, for example by motivating more expressive transform architectures and spatially adaptive variants.
-More broadly, the idea of incorporating signal priors to improve the parameter efficiency of neural codecs may benefit the design of low-complexity neural image and video codecs.
-[…]
-Supplementary Material
-(a) Synthetic-NeRF
-(b) BungeeNeRF
-Figure 5: Comparison of our method with anchor-based methods on two additional datasets, Synthetic-NeRF [mildenhall2021nerf] and BungeeNeRF [xiangli2022bungeenerf].
-Figure 6: Visual comparison between HAC++ and our SHTC method on four scenes: ‘playroom’ (DeepBlending), ‘train’ (Tanks & Temples), ‘garden’ and ‘stump’ (Mip-NeRF360).
-7 Overview of the Supplementary Material
-This supplementary document provides additional technical details, extended experiments, and ablation studies that complement the main paper. Its content is organized as follows:
-  • Sec. 8 compares our method with representative anchor-based methods on two additional datasets, Synthetic-NeRF [mildenhall2021nerf] and BungeeNeRF [xiangli2022bungeenerf].
-  • Sec. 9 presents additional comparisons of computational cost between our approach and CAT-3DGS.
-  • Sec. 10 shows visual comparisons.
-  • Sec. 11 presents more detailed ablation studies.
-  • Sec. 12 provides a more detailed discussion of our transform choices.
-  • Sec. 13 presents additional comparisons in the training time.
-  • Sec. 14 provides further comparisons with a set of widely used baselines and recent methods.
-  • Sec. 15 summarizes implementation details.
-    • Sec. 15.1 explains how we compute pixel-wise distortion in the YCbCr color space.
-    • Sec. 15.2 details the selection of the $\lambda$ values used to generate the R-D curves in the main paper.
-  • Sec. 16 discusses the limitations of our approach and outlines directions for future work.
-  • Sec. 17 reviews additional related work to further contextualize our method.
-8 Additional Results on the Synthetic-NeRF and BungeeNeRF Datasets
-Fig.
-5 compares the R-D curves of our method with several anchor-based 3DGS compression methods on the Synthetic-NeRF [mildenhall2021nerf] and BungeeNeRF [xiangli2022bungeenerf] datasets.
-As shown in the figure, our method consistently achieves better R-D performance, attaining a smaller memory footprint at comparable PSNR levels.
-Tab.
-3 further reports the BD-rate gains of our method over three anchor-based baselines.
-Taken together, these additional experimental results further demonstrate the effectiveness of our approach and serve as a supplementary comparison to the experiments reported in the main text.
-Dataset HAC++ CAT-3DGS (40k) HAC
-Synthetic-NeRF -13.45% - -26.33%
-BungeeNeRF -10.04% -39.90% -52.05%
-Table 3: The BD-rate of our method relative to three baselines: HAC++ [chen2025hac++], CAT-3DGS [zhan2025catdgs] and HAC [chen2024hac].
-Negative values indicate bitrate savings (better) compared with each baseline.
-9 Additional Comparisons with CAT-3DGS
-[…]
-Table 4: Evaluating the effectiveness: CAT-3DGS vs.
-SHTC + simple entropy model.
-From left to right, we report four metrics: encoding time, decoding time, and BD-rate relative to the HAC baseline.
-The best result for each metric is highlighted in bold.
-As a supplement to Table 2 in the main paper, Tab.
-4 compares our method with CAT-3DGS.
-The evaluation metrics include encoding time, decoding time, and BD-rate relative to the HAC baseline.
-With HAC as the common baseline, our method achieves a larger reduction in memory footprint while requiring substantially less encoding and decoding time at inference.
-For CAT-3DGS [zhan2025catdgs], we encountered out-of-memory (OOM) errors during rendering when running the official implementation on our hardware, so the rendering FPS could not be measured.
-However, the training time, encoding time, and decoding time can still be measured before this OOM issue occurs.
-Consequently, we omit the rendering-FPS comparison for CAT-3DGS in the table.
-10 Visual Comparison
-We provide visual comparisons between our SHTC approach and HAC++ [chen2025hac++] in Fig.
-6.
+In this paper, we introduce TTC, a new paradigm for 3DGS compression that jointly optimizes the 3DGS representation, entropy model, and analysis–synthesis transform under a unified R–D objective.
+Compared with anchor-based compressors that solely rely on increasingly complex entropy models and PTC methods that cannot co-adapt the transform with 3DGS, TTC provides a more effective route to redundancy removal.
+Within TTC, we propose SHTC, achieving substantial R–D gains and improved inference efficiency with minimal parameter overhead.
+More broadly, our parameter-efficient design may benefit low-complexity neural image and video codec design.
+Impact Statement
+Our method serves purely as a compression tool, which operates solely on existing data to reduce its storage or transmission cost.
+It neither generates new data nor edits the original content, thereby minimizing risks of misuse, such as data fabrication or manipulation.
+Given its functional scope, we do not anticipate any significant negative societal impacts associated with the deployment of our method.
+On the positive side, by significantly reducing the file size of 3DGS representations and lowering decoding latency, our method facilitates faster downloading and decoding.
+This improves user experience in bandwidth-constrained or real-time applications such as virtual reality, immersive gaming, architectural visualization, and cultural heritage preservation.
+Appendix A Visual Comparison
+We provide visual comparisons between our SHTC approach and HAC++ (Chen et al., 2025c) in Figure 9.
 For each example, we report the bitstream size (in MB) and the PSNR (in dB) of the shown crop.
-To illustrate how SHTC reduces rendering distortion under a smaller or comparable memory footprint, we evaluate four representative scenes: ‘playroom’ (DeepBlending), ‘train’ (Tanks & Temples), and ‘garden’ and ‘stump’ (Mip-NeRF360), which are shown from the first to the fourth row in Fig.
-6.
+To illustrate how SHTC reduces rendering distortion under a smaller or comparable memory footprint, we evaluate four representative scenes: ‘playroom’ (DeepBlending), ‘train’ (Tanks & Temples), and ‘garden’ and ‘stump’ (Mip-NeRF360), which are shown from the first to the fourth row in Figure 9.
 In the ‘playroom’ scene, our method better preserves the structure of the wall-mounted switch, which appears as an unrecognizable blur in the HAC++ reconstruction.
@@
 Figure 10: Ablation Study of SHTC Variants on Mip-NeRF360.
-11 Additional Ablation Study
-11.1 Additional Quantitative Ablation Study
+Appendix B Additional Ablation Study
+B.1 Additional Quantitative Ablation Study on Component Contributions
 In the main paper, we only report the most basic configuration and the complete version of our method, mainly to let readers grasp the overall benefit of SHTC at a glance under the page limit.
@@
 This corresponds to our full model.
-We plot the R–D curves of these variants on the Mip-NeRF360 dataset in Fig.
-7(a).
+We plot the R–D curves of these variants on the Mip-NeRF360 dataset in Figure 10(a).
 Mip-NeRF360 contains diverse and challenging indoor and outdoor scenes, making it a suitable benchmark for ablation with reduced bias.
-From Fig.
-7(a), we observe that introducing a KLT-based base layer for anchor feature compression (HAC-E + BL(F15)) already yields a clear gain over the HAC-E baseline and brings the performance close to HAC++.
+From Figure 10(a), we observe that introducing a KLT-based base layer for anchor feature compression (HAC-E + BL(F15)) already yields a clear gain over the HAC-E baseline and brings the performance close to HAC++.
 Extending the base layer to also compress anchor scalings (HAC-E + BL(F15+S)) and further switching to the YCbCr-based distortion metric (HAC-E + BL(F15+S) + YCbCr) provide additional, consistent improvements.
 The complete SHTC configuration trained with the YCbCr-based distortion (HAC-E + Full SHTC + YCbCr) achieves the best overall R-D trade-off among all variants.
-Tab.
-5 quantifies these effects by reporting BD-rate values over HAC-E on Mip-NeRF360.
+Table 3 quantifies these effects by reporting BD-rate values over HAC-E on Mip-NeRF360.
 HAC++ is also included to facilitate a direct comparison between the two paradigms: using learned transforms with a simple entropy model versus relying solely on a highly complex entropy model.
 The results show that even the simplest transform-based variant (HAC-E + BL(F15)) already surpasses HAC++, and each additional component yields further BD-rate reductions, leading to a substantial $29.33\%$ reduction for HAC-E + Full SHTC + YCbCr.
+Table 3: BD-rate over HAC-E on the Mip-NeRF360 dataset for HAC++ and the proposed SHTC variants.
+Negative values indicate rate savings over HAC-E, and more negative values correspond to better R-D performance.
 BD-rate over HAC-E
@@
 HAC-E+Full SHTC+YCbCr -29.33%
-Table 5: BD-rate over HAC-E on the Mip-NeRF360 dataset for HAC++ and the proposed SHTC variants.
-Negative values indicate rate savings over HAC-E, and more negative values correspond to better R-D performance.
-11.2 Comparison with No-Truncation Variants
+B.2 Comparison with No-Truncation Variants
 Finally, we investigate whether directly transmitting all KLT coefficients can serve as a competitive alternative to our truncated-plus-refinement design.
-In Fig.
-7(b), we first consider a variant that does not truncate the KLT coefficients of anchor features, denoted by HAC-E + BL(F50+S) + YCbCr.
+In Figure 10(b), we first consider a variant that does not truncate the KLT coefficients of anchor features, denoted by HAC-E + BL(F50+S) + YCbCr.
 This configuration directly quantizes and entropy-codes all 50 KLT coefficients.
@@
 Building on HAC-E + BL(F50+S) + YCbCr, we then introduce a monotonically increasing, channel-wise quantization schedule for the 50-dimensional KLT coefficients,
-$q_{s}^{(i)}=q_{s}\exp(\alpha i),\qquad i\in\{0,\ldots,49\},$ (10)
+$q_{s}^{(i)}=q_{s}\exp(\alpha i),\quad i\in\{0,\ldots,49\},$ (8)
 where $\alpha$ is a learnable scalar and $q_{s}$ is the per-anchor base quantization step predicted by HAC.
@@
 (i–j) Corresponding normalized energy distributions for Haar L and Haar H sub-bands.
-12 Discussion of Transform Choices
-12.1 Transform Choices for the Base Layer
+Appendix C Discussion of Transform Choices
+C.1 Transform Choices for the Base Layer
 In this section, we present a high-level comparison of transform design choices, supported by several analysis experiments.
@@
 To guide this choice, we compared their ability to decorrelate anchor channels and compact energy.
-Fig.
-8 shows the absolute inter-channel correlation matrices and the normalized energy distributions of the original HAC features, the DCT coefficients, the Haar coefficients, and the KLT coefficients.
+Figure 11 shows the absolute inter-channel correlation matrices and the normalized energy distributions of the original HAC features, the DCT coefficients, the Haar coefficients, and the KLT coefficients.
 DCT and Haar can reduce inter-channel correlation and slightly improve energy compactness, but their gains are limited: the correlation matrices still exhibit noticeable off-diagonal structure and the energy remains relatively spread out across channels.
@@
 Both networks are 5-layer MLPs, with a total of 21,930 parameters that must be transmitted as part of the bitstream.
-This comparison, evaluated on the Mip-NeRF360 dataset and shown in Fig.
-9, reveals that the black-box MLP transform performs significantly worse in terms of the R-D trade-off than the simple KLT-based design.
+This comparison, evaluated on the Mip-NeRF360 dataset and shown in Figure 12(a), reveals that the black-box MLP transform performs significantly worse in terms of the R-D trade-off than the simple KLT-based design.
 A likely reason is that, under a strict parameter budget, the MLP-based transform lacks sufficient expressive power to accurately reconstruct the input, leading to larger reconstruction errors and thus worse R-D performance.
@@
 On top of this, we use a neural residual coder to compensate for the truncation loss of the KLT coefficients, which further improves the R-D performance over using KLT alone.
-Figure 9: Comparison between ‘HAC-E + KLT (F15)’ and ‘HAC-E + MLP’ on the Mip-NeRF360 dataset.
-12.2 Transform Choices for the Refinement Layer
+(a) Comparison between ‘HAC-E + KLT (F15)’ and ‘HAC-E + MLP’ on the Mip-NeRF360 dataset.
+(b) Offline comparison of our sparsity-aware refinement layer and a black-box MLP-based transform on the bicycle scene of the Mip-NeRF360 dataset.
+The plot shows the $\ell_{1}$ reconstruction loss of the KLT residual versus training iteration.
+Figure 12: Comparison of different design choices.
+C.2 Transform Choices for the Refinement Layer
 When designing the neural refinement layer, a natural question arises: why not simply use a generic MLP?
@@
 We train both models for 3,000 iterations and measure the $\ell_{1}$ reconstruction loss of the KLT residual.
-As shown in Fig.
-10, the sparsity-aware transform consistently drives the loss down to a much lower level than the MLP-based transform, despite using roughly four times fewer parameters.
+As shown in Figure 12(b), the sparsity-aware transform consistently drives the loss down to a much lower level than the MLP-based transform, despite using roughly four times fewer parameters.
 This early-stage verification supports our motivation: leveraging the sparsity prior and using a compressed-sensing-inspired architecture for residual coding yields a highly parameter-efficient refinement module.
 For this reason, we adopt the sparsity-aware design as our default choice for the neural refinement layer in SHTC.
-Figure 10: Offline comparison of our sparsity-aware refinement layer and a black-box MLP-based transform on the bicycle scene of the Mip-NeRF360 dataset.
-The plot shows the $\ell_{1}$ reconstruction loss of the KLT residual versus training iteration.
-13 Training time
-Tab.
-6 compares the training time of our method with three advanced anchor-based 3DGS compression methods that rely on complex entropy models.
-Although our method is not the fastest to train, its training time remains moderate: it requires more training time than HAC++, but consistently less than CAT-3DGS and ContextGS across all three benchmarks.
-Designing a compression system that simultaneously attains the best possible training efficiency, R-D performance, and inference speed is notoriously difficult; in practice, there is no free lunch.
-We argue that the resulting overhead in training time is acceptable because 3DGS compression follows a “compress once, decompress many times” paradigm.
-Training is a one-time offline cost borne by 3DGS asset producers, whereas decoding is performed repeatedly on the user side.
-End users are therefore insensitive to training time but directly affected by download and decoding latency.
-In this context, trading a modest increase in training time for improved R-D performance and faster decoding is both realistic and, in practice, a favorable design choice.
-Our method realizes this trade-off in practice, achieving improved R-D performance and lower decoding latency while keeping the training time within a reasonable range.
-CAT-3DGS ContextGS HAC++ Ours
-Mip-NeRF360 5725 3927 2735 3218
-Tank&Temples 4034 2505 1880 2315
-DeepBlending 3353 2320 1857 2270
-Table 6: Training time (in seconds) of CAT-3DGS, ContextGS, HAC++, and our method on three datasets. The best and second-best results are highlighted in red and yellow cells, respectively.
-Datasets Mip-NeRF360 Tank&Temples DeepBlending
-[…]
-3DGS [kerbl20233d] 27.46 0.812 0.222 750.9 23.69 0.844 0.178 431.0 29.42 0.899 0.247 663.9
-Scaffold-GS [lu2024scaffold] 27.50 0.806 0.252 253.9 23.96 0.853 0.177 86.50 30.21 0.906 0.254 66.00
-Compact3DGS [lee2024compact] 27.08 0.798 0.247 48.80 23.32 0.831 0.201 39.43 29.79 0.901 0.258 43.21
-Compressed3D [niedermayr2024compressed] 26.98 0.801 0.238 28.80 23.32 0.832 0.194 17.28 29.38 0.898 0.253 25.30
-EAGLES [girish2024eagles] 27.14 0.809 0.231 58.91 23.28 0.835 0.203 28.99 29.72 0.906 0.249 52.34
-LightGaussian [fan2024lightgaussian] 27.00 0.799 0.249 44.54 22.83 0.822 0.242 22.43 27.01 0.872 0.308 33.94
-Navaneet et al. [navaneet2024compgs] 27.12 0.806 0.240 19.33 23.44 0.838 0.198 12.50 29.90 0.907 0.251 13.50
-Reduced3DGS [papantonakis2024reducing] 27.19 0.807 0.230 29.54 23.57 0.840 0.188 14.00 29.63 0.902 0.249 18.00
-RDOGaussian [wang2024end] 27.05 0.802 0.239 23.46 23.34 0.835 0.195 12.03 29.63 0.902 0.252 18.00
-PUP 3D-GS [hanson2025pup] 26.67 0.786 0.272 74.65 22.72 0.801 0.244 43.33 28.85 0.881 0.302 69.92
-SizeGS [xie2025sizegs] 27.48 0.806 0.240 18.17 24.04 0.840 0.200 10.93 30.24 0.903 0.271 7.92
-FlexGaussian [tian2025flexgaussian] 26.38 0.780 0.251 40.80 22.44 0.804 0.219 16.30 28.61 0.884 0.269 25.48
-NeuralGS [tang2025neuralgs] 27.35 0.806 0.240 16.90 23.63 0.841 0.192 12.06 29.91 0.906 0.254 12.98
-LocoGS [shin2025localityaware] 27.40 0.815 0.219 13.89 23.89 0.854 0.160 12.34 30.17 0.906 0.244 13.39
-FCGS (low rate) [chen2025fast] 27.05 0.798 0.237 36.30 23.48 0.833 0.193 18.80 29.27 0.893 0.257 30.10
-FCGS (high rate) [chen2025fast] 27.39 0.806 0.226 67.20 23.62 0.839 0.184 33.60 29.58 0.899 0.248 54.50
-SOG [morgenstern2024compact] 26.56 0.791 0.241 16.70 23.15 0.828 0.198 9.30 29.12 0.892 0.270 5.70
-CodecGS [lee2025compression] 27.30 0.810 0.236 9.78 23.63 0.842 0.192 7.46 29.82 0.907 0.251 8.62
-MesonGS [xie2024mesongs] 26.99 0.796 0.247 27.16 23.32 0.837 0.193 16.99 29.51 0.901 0.251 24.76
-CompGS [liu2024compgs] 27.26 0.803 0.239 16.50 23.70 0.837 0.208 9.60 29.69 0.901 0.279 8.77
-FP-Net [ma2025enhancing] 27.46 0.801 0.249 11.08 24.13 0.847 0.192 6.34 30.14 0.906 0.274 3.46
-FP-GS [tang2025feature] 27.53 0.806 0.240 12.82 24.32 0.811 0.185 7.32 30.05 0.906 0.266 3.91
-HAC [chen2024hac] 27.53 0.807 0.238 15.26 24.04 0.846 0.187 8.10 29.98 0.902 0.269 4.35
-Context-GS [wang2024contextgs] 27.62 0.808 0.237 12.68 24.20 0.852 0.184 7.05 30.11 0.907 0.258 3.45
-HAC++ [chen2025hac++] 27.60 0.803 0.253 8.34 24.22 0.849 0.190 5.18 30.16 0.907 0.266 2.91
-Ours 27.69 0.804 0.248 8.10 24.34 0.845 0.191 5.19 30.19 0.903 0.272 2.49
 Table 4: Comparison of our SHTC framework with other 3DGS data compression methods, including 3DGS and Scaffold-GS, for reference.
@@
 The size values are measured in megabytes (MB).
-14 Additional Comparisons with Other Methods
+[…]
+(a) Mip-NeRF360
+(b) Tank&Temples
+(c) DeepBlending
+Figure 13: Comparison of our method with several conceptually representative baslines.
+Appendix D Additional Comparisons with Other Methods
 In addition to the comparisons with anchor-based compression methods using R–D curves and BD-rate, we also evaluate our method against a broad set of anchor-free and hybrid baselines.
-These include several commonly used 3DGS and compression baselines [kerbl20233d, lu2024scaffold, lee2024compact, niedermayr2024compressed, girish2024eagles, fan2024lightgaussian, morgenstern2024compact, navaneet2024compgs, papantonakis2024reducing, wang2024end, liu2024compgs], the feed-forward compression method FCGS [chen2025fast], and two neural-field-based 3DGS compression methods, LocoGS [shin2025localityaware] and NeuralGS [tang2025neuralgs].
-We further consider several recent 3DGS compression methods [hanson2025pup, xie2025sizegs, tian2025flexgaussian, lee2025optimized], two HAC-based approaches that introduce prediction modules to improve HAC [tang2025feature, ma2025enhancing], and a group of methods that apply transform coding in a post-training stage for 3DGS compression [morgenstern2024compact, lee2025compression, xie2024mesongs, yang2025hybridgs].
-Regarding SOGS [zhang2025sogs], we do not include it in our quantitative comparison for two reasons.
+These include several commonly used 3DGS and compression baselines (Kerbl et al., 2023; Lu et al., 2024; Lee et al., 2024; Niedermayr et al., 2024; Girish et al., 2024; Fan et al., 2024; Morgenstern et al., 2024; Navaneet et al., 2024; Papantonakis et al., 2024; Wang et al., 2024a; Liu et al., 2024b), the feed-forward compression method FCGS (Chen et al., 2025b), and two neural-field-based 3DGS compression methods, LocoGS (Shin et al., 2025) and NeuralGS (Tang et al., 2025b).
+We further consider several recent 3DGS compression methods (Hanson et al., 2025; Xie et al., 2025; Tian et al., 2025; Lee et al., 2025a), two HAC-based approaches that introduce prediction modules to improve HAC (Tang et al., 2025a; Ma et al., 2025), and a group of methods that apply transform coding in a post-training stage for 3DGS compression (Morgenstern et al., 2024; Lee et al., 2025b; Xie et al., 2024; Yang et al., 2025).
+Regarding SOGS (Zhang et al., 2025), we do not include it in our quantitative comparison for two reasons.
 First, to the best of our knowledge, there is no official implementation available, and the original paper does not report file sizes or distortion metrics on the datasets used in our experiments, which makes a fair and reproducible comparison difficult.
 Second, SOGS is essentially a 3DGS compaction method rather than a full compression scheme with an explicit bitstream and entropy coding.
-Based on the model statistics reported in [zhang2025sogs], we can only roughly infer an effective compression ratio of about $2\times$ over Scaffold-GS; this is merely a rough estimate, since the authors do not provide explicit compression ratios.
+Based on the model statistics reported in (Zhang et al., 2025), we can only roughly infer an effective compression ratio of about $2\times$ over Scaffold-GS; this is merely a rough estimate, since the authors do not provide explicit compression ratios.
 In contrast, by employing the proposed SHTC framework, our method achieves substantially higher compression ratios over Scaffold-GS.
@@
 Because many of the above methods report only one or two operating points on the R–D curve, a meaningful BD-rate computation is not feasible.
-For these methods, we therefore provide a numerical comparison in tabular form in Tab.
-7.
-As shown in Tab.
-7, our method achieves a better memory-quality trade-off than above approaches.
+For these methods, we therefore provide a numerical comparison in tabular form in Table 4.
+As shown in Table 4, our method achieves a better memory-quality trade-off than above approaches.
+[…]
+Table 5: Summary of $\lambda$ values used to generate R-D curves on different datasets.
 HAC ContextGS HAC++ Ours Mip-NeRF360 $\{0.002,0.004,0.008,0.015\}$ $\{0.002,0.004,0.008,0.015,0.025\}$ $\{0.002,0.004,0.008,0.015,0.025\}$ $\{0.002,0.004,0.008,0.015\}$ Tank&Temples $\{0.002,0.004,0.008,0.015,0.025\}$ $\{0.002,0.004,0.008,0.015,0.025,0.04\}$ $\{0.002,0.004,0.008,0.015,0.025\}$ $\{0.002,0.004,0.008,0.015\}$ DeepBlending $\{0.002,0.004,0.008,0.015,0.025\}$ $\{0.002,0.004,0.008,0.015,0.025,0.04\}$ $\{0.002,0.004,0.008,0.015,0.025,0.04\}$ $\{0.002,0.004,0.008,0.015,0.02,0.025\}$ Synthetic-NeRF $\{0.0005,0.001,0.002,0.003,0.004\}$ - $\{0.0005,0.001,0.002,0.003,0.004\}$ $\{0.0005,0.001,0.002,0.003,0.004\}$ BungeeNeRF $\{0.001,0.002,0.003,0.004,0.006\}$ $\{0.001,0.004\}$ $\{0.0005,0.001,0.002,0.003,0.004,0.006,0.008\}$ $\{0.001,0.002,0.003,0.004,0.006\}$
-Table 8: Summary of $\lambda$ values used to generate R-D curves on different datasets.
-15 Implementation Details
-15.1 Pixel-wise Error in YCbCr Color Space
+Appendix E Implementation Details
+The overall training loss is defined as
+$\mathcal{L}=\mathcal{L}_{d}+\lambda\mathcal{L}_{\mathrm{b-rate}}+\lambda_{o}\mathcal{L}_{\mathrm{reg}}+\lambda_{e}\ell_{1}(\mathbf{r},\hat{\mathbf{r}})+\lambda_{r}\mathcal{L}_{\mathrm{r-rate}}$ (9)
+[…]
+E.1 Pixel-wise Error in YCbCr Color Space
 We propose a new distortion term for supervising the rendered images, inspired by the YCbCr color transform used in traditional image and video codecs.
@@
 Together, the distortion term is defined as
-$\mathcal{L}_{\mathrm{YCbCr}}=$ $\ell_{1}(I_{Y},\hat{I}_{Y})+\lambda_{c}\ell_{1}(I_{Cb},\hat{I}_{Cb})+\lambda_{c}\ell_{1}(I_{Cr},\hat{I}_{Cr})$
-$+$ $\lambda_{H}\ell_{1}(\nabla^{2}I_{Y},\nabla^{2}\hat{I}_{Y})$ (11)
-$+$ $\lambda_{TV}\mathrm{TV}(\hat{I}_{Cb})+\lambda_{TV}\mathrm{TV}(\hat{I}_{Cr}).$
+$\mathcal{L}_{\mathrm{YCbCr}}=\ell_{1}(I_{Y},\hat{I}_{Y})+\lambda_{c}\ell_{1}(I_{Cb},\hat{I}_{Cb})+\lambda_{c}\ell_{1}(I_{Cr},\hat{I}_{Cr})+\lambda_{H}\ell_{1}(\nabla^{2}I_{Y},\nabla^{2}\hat{I}_{Y})+\lambda_{TV}\mathrm{TV}(\hat{I}_{Cb})+\lambda_{TV}\mathrm{TV}(\hat{I}_{Cr})$ (10)
 This distortion term acts as a regularizer on the compressed 3DGS representation, encouraging it to render images with reduced luminance errors at a given bitrate, since luminance errors are perceptually more important than chrominance errors.
@@
 In our implementation, we set $\lambda_{c}$, $\lambda_{H}$, and $\lambda_{TV}$ to 0.6, 0.15, and 0.1, respectively.
-15.2 Details of $\lambda$ Selection for R-D Curves
-In this subsection, we summarize in Tab.
-8 the $\lambda$ values used to generate the R–D curves in our experiments.
-In addition to these values, we also clarify several special cases regarding how $\lambda$ and the corresponding R-D points are obtained for some baselines:
-  • Main text.
-For the R-D curves in the main paper, we train the model with $\lambda\in\{0.002,0.004,0.008,0.015\}$, which yields a PSNR range of roughly 1 dB.
-The spacing between consecutive $\lambda$ value is gradually increased so as to cover a wide bitrate range while providing approximately uniform sampling of the R–D curves within that range.
-Because different models and datasets exhibit different sensitivity to $\lambda$, we additionally use larger $\lambda$ values for some baselines so that the PSNR range of each curve is comparable to that of the others.
+E.2 Details of $\lambda$ Selection for R-D Curves
+In this subsection, we summarize the $\lambda$ values used to obtain the R-D curves in our experiments in Table 5.
+For our primary experiments, we set the Lagrange multiplier $\lambda$ to $\{0.002,0.004,0.008,0.015\}$, which yields operating points spanning a wide rate-distortion range and enables a comprehensive comparison.
+To obtain more uniformly spaced operating points along the R–D curve, we use progressively larger gaps between successive $\lambda$ values.
+This is motivated by the highly nonlinear relationship between $\lambda$ and the achieved rate.
+In particular, in the low-rate regime, larger changes in $\lambda$ are often necessary to produce clearly separated rates.
+In addition, we discuss several special cases for baselines, detailing how $\lambda$ (or its equivalent control parameter) and the corresponding R–D points are obtained.
   • CAT-3DGS.
-For CAT-3DGS [zhan2025catdgs], we encountered out-of-memory (OOM) errors during rendering when using the official implementation on our hardware.
+For CAT-3DGS (Zhan et al., 2025), we encountered out-of-memory (OOM) errors during rendering when using the official implementation on our hardware.
 Consequently, we do not re-train or re-evaluate this method; instead, we directly use the R-D points reported in the original paper as its results in our plots.
@@
   • PCGS.
-For PCGS, on both Synthetic-NeRF and BungeeNeRF, we found it difficult to obtain a sufficiently wide PSNR range for a stable BD-rate estimation by tuning $\lambda$.
-Therefore, we only plot the R-D points reported in the original paper and treat PCGS as a qualitative reference rather than including it in the quantitative BD-rate comparison.
-16 Limitation and Future Work
+For PCGS (Chen et al., 2025a), on DeepBlending, Mip-NeRF360, and Tanks&Temples, we train the model with $\lambda\in\{0.004,0.008,0.015,0.025\}$ to enlarge the distortion range.
+Nevertheless, even with this wider distortion span, we still cannot obtain sufficient PSNR overlap with other methods, and we find it generally difficult to further increase the overlap for a meaningful BD-rate computation.
+Therefore, we include these PCGS results only for plotting the R-D curves as a qualitative comparison, and we do not include PCGS in the quantitative BD-rate evaluation.
+Given that expanding the overlap is consistently challenging on the above three datasets, for Synthetic-NeRF and BungeeNeRF we do not introduce new $\lambda$ values for re-training; instead, we directly use the results reported by the original authors as a reference in the R-D plots (also excluded from BD-rate).
+We note that our evaluation setup is largely comparable to PCGS, since both are derived from the HAC/HAC++ codebase.
+Appendix F Limitation and Future Work
 In this work, SHTC adopts transforms that is shared across all spatial regions.
@@
 As our primary goal in this work is to demonstrate the necessity and effectiveness of incorporating transform coding into the 3DGS training process, we leave the design and evaluation of these spatially adaptive variants for future work.
-17 Supplementary Related Work
-17.1 Unstructured 3DGS compression methods
+Appendix G Supplementary Related Work
+G.1 Unstructured Compression Paradigm
 3DGS has recently emgered as a pioneering approach for 3D reconstruction and representation, offering both high-quality and real-time rendering.
@@
 The challenge lies in identifying these unimportant Gaussians.
-[navaneet2024compgs] simply uses opacity as a criterion to remove transparent or nearly invisible Gaussians.
-[ali2024elmgs, ali2024trimming] further use both gradient magnitude and opacity levels as reference to identify removable Gaussians.
-Several subsequent studies assess each Gaussian’s contribution to ray color prediction and utilize this as the importance score [fan2024lightgaussian, girish2024eagles, niemeyer2024radsplat].
+(Navaneet et al., 2024) simply uses opacity as a criterion to remove transparent or nearly invisible Gaussians.
+(Ali et al., 2024a, b) further use both gradient magnitude and opacity levels as reference to identify removable Gaussians.
+Several subsequent studies assess each Gaussian’s contribution to ray color prediction and utilize this as the importance score (Fan et al., 2024; Girish et al., 2024; Niemeyer et al., 2024).
 Lee et al.
-learn a binary mask to prune those unimportant Gaussians [lee2024compact], which is adopted by several subsequent works [wang2024end].
-[hanson2025pup] propose a principled sensitivity pruning score by computing a second-order approximation of the reconstruction error with respect to the parameters of each Gaussian.
+learn a binary mask to prune those unimportant Gaussians (Lee et al., 2024), which is adopted by several subsequent works (Wang et al., 2024a).
+(Hanson et al., 2025) propose a principled sensitivity pruning score by computing a second-order approximation of the reconstruction error with respect to the parameters of each Gaussian.
 SH distillation
@@
 This leads to significant waste in using high-degree SHs.
-To address this, [papantonakis2024reducing] assign each Gaussian an appropriate SH degree to reduce waste, while [fan2024lightgaussian] distill knowledge from a high-degree SH teacher model to a student model with truncated, lower-degree SHs.
+To address this, (Papantonakis et al., 2024) assign each Gaussian an appropriate SH degree to reduce waste, while (Fan et al., 2024) distill knowledge from a high-degree SH teacher model to a student model with truncated, lower-degree SHs.
 Scalar quantization
@@
 However, opacity, scale, rotation, and SH coefficients are tolerant to small inaccuracies, suggesting that lower bit-depths can represent them with minimal performance drop.
-In light of this, [papantonakis2024reducing] propose a scalar quantization scheme that assigns each full-precision value to its nearest codeword in a K-means-derived codebook, storing only the corresponding index.
-[girish2024eagles] learn a latent representation, quantize it using a uniform scalar quantizer, and employ a decoder to predict Gaussian attributes from the quantized representation.
+In light of this, (Papantonakis et al., 2024) propose a scalar quantization scheme that assigns each full-precision value to its nearest codeword in a K-means-derived codebook, storing only the corresponding index.
+(Girish et al., 2024) learn a latent representation, quantize it using a uniform scalar quantizer, and employ a decoder to predict Gaussian attributes from the quantized representation.
 Vector quantization
@@
 Instead of storing high-precision vectors, only compact indices are needed.
-Based on this insight, several works apply vector quantization to eliminate the redundancy within Gaussian attributes [navaneet2024compgs, fan2024lightgaussian, niedermayr2024compressed].
-Furthermore, [wang2024end] employ entropy-constrained vector quantization, incorporating entropy coding into the quantization process.
+Based on this insight, several works apply vector quantization to eliminate the redundancy within Gaussian attributes (Navaneet et al., 2024; Fan et al., 2024; Niedermayr et al., 2024).
+Furthermore, (Wang et al., 2024a) employ entropy-constrained vector quantization, incorporating entropy coding into the quantization process.
 It assigns shorter codes to those code vectors that are used more frequently, effectively compressing the index stream.
-17.2 Transform coding in NeRF/3DGS compression
-Before the emergence of 3DGS, grid-based NeRF representations [muller2022instant, fridovich2023k, sun2022direct, chen2022tensorf] were considered a promising alternative to the classic NeRF [mildenhall2021nerf].
-Several works have explored the use of transform coding to compress the feature grids in these representations [rho2023masked, li2024nerfcodec, pham2024neural].
-Specifically, Rho et al.
-use a Discrete Wavelet Transform (DWT) to exploit transform-domain sparsity for compressing feature planes [rho2023masked].
-Li et al.
-repurpose neural image codecs as learned nonlinear transforms: for each scene, they pre-train a plane-based NeRF, then scene-specifically fine-tune a subset of neural codec layers and encode the pretrained feature planes with the fine-tuned codec [li2024nerfcodec].
-Pham et al.
-follow the same two-stage recipe but replace the off-the-shelf codec with lightweight nonlinear transforms designed from scratch [pham2024neural].
-Compared with grid-based NeRFs, applying transform coding to 3DGS is more challenging because 3DGS data are irregular and unordered.
+G.2 Post-training Transform Coding Paradigm
+In this subsection, we focus on methods that employ transforms during the post-training stage of 3DGS compression.
+Applying transform coding to 3DGS is challenging because 3DGS data are irregular and unordered.
 To sidestep these challenges, prior work defers the application of transform coding until the post-training stage, once the 3DGS geometry is fixed.
 Two main strategies have been explored.
-First, some methods project 3DGS onto a set of feature planes [morgenstern2024compact] or adopt a triplane representation [chan2022efficient] to implicitly encode Gaussian attributes [lee2025compression], then compress the learned planes with standard image or video codecs.
-Second, other methods treat a pretrained 3DGS model as a point cloud and apply point-cloud attribute transforms such as Graph Fourier transform (GFT) [sandryhaila2013discrete] and Region-adaptive Hierarchical Transform (RAHT) [de2016compression] to compress the model [xie2024mesongs, yang2024benchmark, huang2025hierarchical, wang2025adaptive].
+First, some methods project 3DGS onto a set of feature planes or adopt a triplane representation (Chan et al., 2022) to implicitly encode Gaussian attributes, then compress the learned planes with standard image or video codecs (Lee et al., 2025b; Morgenstern et al., 2024).
+Second, other methods treat a pretrained 3DGS model as a point cloud and apply point-cloud attribute transforms such as Graph Fourier transform (GFT) (Sandryhaila and Moura, 2013) and Region-adaptive Hierarchical Transform (RAHT) (De Queiroz and Chou, 2016) to compress the model (Xie et al., 2024; Yang et al., 2024; Huang et al., 2025; Wang et al., 2025; Yang et al., 2025).
 The above methods borrow transforms from images, video, or point clouds but apply fixed, linear transforms only after training.
 This decoupling blocks mutual adaptation between the transform and the 3DGS representation, often limiting R–D performance.
-In contrast, we propose an end-to-end transform-coding framework that jointly adapts the transform and the 3DGS representation, improving compression performance over post-training transform approaches.
+In contrast, we propose an end-to-end transform-coding framework that jointly adapts the transform, entropy model and the 3DGS representation, significantly improving compression performance over post-training transform approaches.